### v1

# Skill-aware Mutual Information Optimisation for Generalisation in Reinforcement Learning

Abstract
Meta-Reinforcement Learning (Meta-RL) agents can struggle to operate across tasks with varying environmental features that require different optimal skills (i.e., different modes of behaviours). Using context encoders based on contrastive learning to enhance the generalisability of Meta-RL agents is now widely studied but faces challenges such as the requirement for a large sample size, also referred to as the $\log$-$K$ curse. To improve RL generalisation to different tasks, we first introduce Skill-aware Mutual Information (SaMI), an optimisation objective that aids in distinguishing context embeddings according to skills, thereby equipping RL agents with the ability to identify and execute different skills across tasks. We then propose Skill-aware Noise Contrastive Estimation (SaNCE), a $K$-sample estimator used to optimise the SaMI objective. We provide a framework for equipping an RL agent with SaNCE in practice and conduct experimental validation on modified MuJoCo and Panda-gym benchmarks. We empirically find that RL agents that learn by maximising SaMI achieve substantially improved zero-shot generalisation to unseen tasks. Additionally, the context encoder equipped with SaNCE demonstrates greater robustness to reductions in the number of available samples, thus possessing the potential to overcome the $\log$-$K$ curse.
1 Introduction
Meta-Reinforcement Learning (Meta-RL) agents often learn policies that do not generalise across tasks in which the environmental features and optimal skills are different [des Combes et al., 2018, Garcin et al., 2024]. Consider, for example, a set of cube-moving tasks, in which an agent is required to move a cube to a goal position on a table (Figure 1). These tasks become challenging if the underlying environmental features vary between different tasks, such as table friction. Suppose an agent
Figure 1: (a) In a cube-moving environment, tasks are defined according to different environmental features. (b) Different tasks have different transition dynamics caused by underlying environmental features, hence optimal skills are different across tasks.
attempts to push a cube across a table covered by a tablecloth and receives environmental feedback showing that the cube has become "unpushable". In this scenario, a generalist agent should infer that the table friction is relatively high. Consequently, the agent ought to adapt its strategy, opting to lift the cube off the table to avoid friction, rather than continuing to push. Recent advances in Meta-RL [Lee et al., 2020, Agarwal et al., 2021, Mu et al., 2022, Dunion et al., 2023b, a] have shown promising potential to perceive and understand environmental features by inferring context embeddings from a small number of interactions in the environment. The context embedding is expected to capture the distribution of tasks and efficiently infer new tasks. Meta-RL methods then train a policy conditioned on the context embedding to generalise to multiple tasks.
The context encoder is the key component for capturing the context embedding from recent experiences [Clavera et al., 2019a, Lee et al., 2020], which affects generalisation performance significantly. Some Meta-RL algorithms [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022] are equipped with context encoders based on contrastive learning, which uses the InfoNCE lower bound [Oord et al., 2019] to optimise the mutual information (MI) between context embeddings. InfoNCE is a $K$-sample MI estimator and is utilised to learn distinct context embeddings for each task. Despite significant advancements, integrating contrastive learning with Meta-RL poses several unresolved challenges, of which two are particularly relevant to this research: (i) existing context encoders based on contrastive learning do not distinguish tasks that require different skills; many prior algorithms only pull embeddings of the same tasks together and push those of different tasks apart. However, for example, a series of cube-moving tasks with high friction may only require a Pick&Place skill (picking the cube off the table and placing it at the goal position), making further differentiation unnecessary. (ii) K-sample MI estimators are sensitive to the sample size K (i.e., the log-K curse) [Poole et al., 2019]; a substantial quantity of negative samples is required to approximate the true MI, which is challenging due to RL’s low sample efficiency [Franke et al., 2021] and often impractical for achieving accurate MI estimation [Arora et al., 2019, Nozawa and Sato, 2021]. The effectiveness of $K$-sample MI estimators breaks down with a small sample size and leads to a significant performance drop in downstream RL tasks [Mnih and Teh, 2012, Guo et al., 2022].
To enhance RL generalisation across different tasks, we propose that the context embeddings should optimise downstream tasks and indicate whether the current skill remains optimal or requires further exploration, thereby addressing issue (i). This approach also reduces the necessary sample size and helps to overcome issue (ii) by concentrating solely on extracting task-relevant MI. Specifically, we propose a three-step process tailored to RL: (1) We introduce Skill-aware Mutual Information (SaMI), a smaller ground-truth MI that discriminates context embeddings according to skills by maximising the MI between context embedding, skills and trajectories. Additionally, we provide a theoretical proof of why introducing skills can make the ground-truth MI smaller, thus making it easier to optimise; (2) We propose a more data-efficient $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), used to optimise SaMI and reduce the negative sample space based on skills to overcome the $log$-$K$ curse; (3) We provide a framework to equip any Meta-RL algorithm with SaNCE in practice, and propose a practical skill-aware sampling method for SaNCE when the skills are not known in advance.
We demonstrate empirically in MuJoCo [Todorov et al., 2012] and Panda-gym [Gallouédec et al., 2021] that SaMI improves the zero-shot generalisation performance in sets of previously unseen tasks with moderate and extreme difficulty. In the MuJoCo benchmark, SaMI enhances two Meta-RL algorithms [Yu et al., 2020, Fu et al., 2021] by achieving an average increase of 42.27% in returns for moderate testing tasks and a 53.14% improvement for extreme testing tasks. Meanwhile, in the Panda-gym benchmark, SaNCE boosts success rates by an average of 20.23% in moderate tests and 18.36% in extreme tests. This indicates SaMI’s advantage in encoding information that enables an agent to execute effective skills in downstream control tasks. SaNCE-based RL algorithms utilise smaller sample spaces while achieving improved downstream control performance, indicating their potential to overcome the $\log$-$K$ curse. Visualisation of the learned context embeddings shows distinct clusters corresponding to different skills, indicating that SaMI is a step towards more robust and versatile RL agents.
Figure 2: A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$. Executing different skills (e.g., Push skill $\pi(\cdot|c_{1})$ and Pick&Place skill $\pi(\cdot|c_{2})$) shows different modes of behaviours in trajectories (e.g., $\tau^{+}_{c_{1}}$ where the cube moves on the table; $\tau^{+}_{c_{2}}$ where the cube leaves the table).
2 Related works
Meta-RL. Meta-RL methods train an agent conditioned on context embeddings to improve generalisation to unseen tasks. As the key component of Meta-RL, the quality of the context embedding can significantly affect the agent’s performance. Existing algorithms can be categorised into three types based on different context embeddings. In the first category, the context embedding is learned by minimising the downstream RL loss [Rakelly et al., 2019, Yu et al., 2020]. PEARL [Rakelly et al., 2019] learns probabilistic context embeddings by recovering the value function. Multi-task SAC + TE (TESAC) [Yu et al., 2020] uses the Task Embedding (TE) [Hausman et al., 2018] as input to policies, parameterising the learned policies via a shared embedding space and maximising average returns. However, the update signals from the RL loss are stochastic and weak, and may not capture the similarity relations among tasks [Fu et al., 2021]. The second category involves learning context embeddings through dynamics prediction [Lee et al., 2020, Zhou et al., 2019], which can make the context embeddings noisy, as they may model irrelevant dependencies and overlook task-specific information [Fu et al., 2021]. The third category employs contrastive learning [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022], achieving significant improvements in context learning. However, these methods overlook the similarity of skills between different tasks, thus failing to achieve effective zero-shot generalisation by executing different skills. Our improvements build upon this third category by distinguishing context embeddings according to different optimal skills.
Contrastive learning. Contrastive learning has been applied to RL due to its significant momentum in representation learning in recent years, attributed to its superior effectiveness [Tishby and Zaslavsky, 2015, Hjelm et al., 2019, Dunion and Albrecht, 2024], ease of implementation [Oord et al., 2019], and strong theoretical connection to mutual information (MI) estimation [Poole et al., 2019]. MI is often estimated through noise-contrastive estimation (NCE) [Gutmann and Hyvärinen, 2010], InfoNCE [Oord et al., 2019], and variational objectives [Hjelm et al., 2019]. InfoNCE has garnered recent interest due to its lower variance [Song and Ermon, 2020] and superior performance in downstream tasks. However, InfoNCE may underestimate true MI because it is constrained by the number of samples $K$. To address this issue, CCM [Fu et al., 2021] leverages InfoNCE to learn the context embedding with a large number of samples. DOMINO [Mu et al., 2022] reduces the total MI by introducing an independence assumption; however, this results in biased information. Correspondingly, we focus on proposing an unbiased alternative MI objective and a more data-efficient $K$-sample estimator tailored for downstream RL tasks, which, to our knowledge, have not been addressed in previous research.
3 Preliminaries
Reinforcement learning. In Meta-RL, we assume an environment is a distribution $\xi(e)$ of tasks $e$ (e.g. uniform in our experiments). Each task $e\sim\xi(e)$ has a similar structure that corresponds to a Markov Decision Process (MDP) [Puterman, 2014], defined by $\mathcal{M}_{e}=(\mathcal{S},\mathcal{A},R,P_{e},\gamma)$, with a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $R(s_{t},a_{t})$ where $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$, state transition dynamics $P_{e}(s_{t+1}|s_{t},a_{t})$, and a discount factor $\gamma\in[0,1)$. In order to address the problem of zero-shot generalisation, we consider the transition dynamics $P_{e}(s_{t+1}|s_{t},a_{t})$ vary across tasks $e\sim\xi(e)$ according to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$ that are not included in states $s$ and can be continuous random variables, such as mass and friction, or discrete random variables, such as the cube’s material. For instance, in a cube-moving environment (Figure 1), an agent has different tasks that are defined by different environmental features (e.g., mass and friction). The Meta-RL agent’s goal is to learn a generalisable policy $\pi$ that is robust to such dynamic changes. Specifically, given a set of training tasks $e$ sampled from $\xi_{\text{train}}(e)$, we aim to learn a policy that can maximise the discounted returns, $\arg\max_{\pi}\mathbb{E}_{e\sim\xi_{\text{train}}(e)}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim P_{e}(s_{t+1}|s_{t},a_{t})]$, and can produce accurate control for unseen test tasks $e$ sampled from $\xi_{\text{test}}(e)$.
Contrastive learning. A Meta-RL agent first consumes a collected trajectory $\tau_{t:t+T}=\{s_{t},a_{t},r_{t},s_{t+1},...,s_{t+T},a_{t+T},r_{t+T}\}$ and outputs a context embedding $c$ through a context encoder $\psi(c|\tau_{t:t+T})$, then executes policy $\pi$ conditioned on the current state $s_{t}$ and context embedding $c$ to work across multiple tasks. As a key component, a context encoder $\psi$ captures the unknown context $c$ of the task. We focus on a context encoder that optimises the InfoNCE bound $I_{\text{InfoNCE}}(x;y)$, which is a $K$-sample estimator of the MI $I(x;y)$ [Oord et al., 2019]. Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, $I_{\text{InfoNCE}}(x;y)$ is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$:
$I_{\text{InfoNCE}}(x;y|\psi,K)=\mathbb{E}\left[\log\frac{f_{\psi}(x,y_{1})}{\frac{1}{K}\sum_{k=1}^{K}f_{\psi}(x,y_{k})}\right].$ (1)
InfoNCE constructs a formal lower bound to the MI, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\leq I(x;y)$ [Guo et al., 2022, Chen et al., 2021]. Given two inputs $x$ and $y$, their embedding similarity is $f_{\psi}(x,y)=e^{\psi(x)^{\top}\cdot\psi(y_{1})/\beta}$, where $\psi$ is the context encoder that projects $x$ and $y$ into the context embedding space, the dot product is used to calculate the similarity score between $\psi(x),\psi(y)$ pairs [Wu et al., 2018, He et al., 2020], and $\beta$ is a temperature hyperparameter that controls the sensitivity of the product. Some previous Meta-RL methods [Lee et al., 2020, Mu et al., 2022] learn a context embedding $c$ by maximising $I_{\text{InfoNCE}}(c;\tau_{c}|\psi,K)$ between the context $c$ embedded from a trajectory in the current task, and the historical trajectories $\tau_{c}$ under the same environmental features setting.
4 Skill-aware mutual information optimisation for Meta-RL
4.1 The $log$-$K$ curse of $K$-sample MI estimators
In this section, we provide a theoretical analysis of the challenge inherent in learning a $K$-sample estimator for MI, commonly referred to as the $log$-$K$ curse. Based on this theoretical analysis, we give insights to overcome this challenge.
If we have sufficient training epochs for the context encoder $\psi$ in Equation 1 and a sample size $K$, $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K$ [Mnih and Teh, 2012, Guo et al., 2022]. Hence, the MI we can optimise is bottlenecked by the number of available samples, formally expressed as:
Lemma 1
[…]
To ensure that the $K$-sample estimator is a tight bound of the ground-truth MI, inspired by Lemma 1, we derive three key insights when learning a context encoder: (1) focus on a ground-truth MI that is smaller than $I(x;y)$ and achievable with a smaller sample size; (2) develop a $K$-sample estimator tighter than $I_{\text{InfoNCE}}$; (3) increasing sample quantity $K$, however, this is usually impractical.
A meta-RL agent learns a context encoder by maximising MI between trajectories $\tau_{c}$ and context embeddings $c$. Driven by insight (1), we introduce Skill-aware Mutual Information (SaMI) in Section 4.2, designed to enhance the zero-shot generalisation of downstream RL tasks. Corresponding to insight (2), we propose Skill-aware Noise Contrastive Estimation (SaNCE) to maximise SaMI with limited samples in Section 4.3. Finally, Section 4.4 demonstrates how to equip any Meta-RL agent with SaNCE in practice.
Figure 3: A $K$-sample estimator $I_{\text{InfoNCE}(c;\pi_{c};\tau_{c})}$, with limited sample size $K$, is a loose lower bound of $I(c;\tau_{c})$ and leads to lower performance embeddings. $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ is a lower ground-truth MI, and $I_{\text{SaNCE}}(c;\pi_{c};\tau_{c})$ is a tighter lower bound of $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$.
4.2 Skill-aware mutual information: a smaller ground-truth MI
An useful tool in learning a versatile agent is to understand when to explore novel skills or switch between existing skills in multi-task settings. To start with, we define skills:
Definition 1 (Skills)
A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$, abbreviated as $\pi_{c}$. If a skill $\pi_{c}$ is conditioned on a state $s_{t}$, we can sample actions $a_{t}\sim\pi(\cdot|c,s_{t})$. If we sample actions from $\pi_{c}$ at consecutive timesteps, we gain a trajectory $\tau_{c}=\{s_{t},a_{t},r_{t},s_{t+1},...,s_{t+T},a_{t+T},r_{t+T}\}$.
After executing skills, an agent should distinguish tasks through environmental feedback to be aware of underlying environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. In this regard, the context encoder $\psi$ should learn by maximising the MI between context embedding $c$, skills $\pi_{c}$, and environmental feedback (i.e., trajectories $\tau_{c}$). We achieve this learning process by maximising the MI $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$, in which we introduce a variable, skill $\pi_{c}$, into $I(c;\tau_{c})$. Formally, we propose a novel MI optimisation objective for a context encoder, Skill-aware Mutual Information (SaMI), which is defined as:
$I_{\text{SaMI}}(c;\pi_{c};\tau_{c})=\mathbb{E}_{p(c,\pi_{c},\tau_{c})}\left\{\log\ \frac{p(\tau_{c}|c,\pi_{c})}{p(\tau_{c})}\right\}.$ (2)
Although we cannot evaluate $p(c,\pi_{c},\tau_{c})$ directly, we approximate it by Monte-Carlo sampling, using $K$ samples from $p(c,\pi_{c},\tau_{c})$. After introducing the variable $\pi_{c}$, $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ contains less information, i.e., $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$ (see proof in Appendix B). So the context encoder $\psi$ has a lower ground-truth MI $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ to approach. $I_{\text{SaMI}}$ also enables the RL agent to embody diverse skills and switch skills based on the context embedding $c$. In other words, the skill $\pi_{c}$ conditioned on different context embeddings $c$ will exhibit various modes of behaviour, such as the Push skill $\pi_{c_{1}}$ (pushing a cube on the table to the target position) or the Pick&Place skill $\pi_{c_{2}}$.
Figure 4: A comparison of sample spaces for task $e_{1}$. Positive samples $\tau_{c_{1}}$ or $\tau_{c_{1}}^{+}$ are always from current task $e_{1}$. For SaNCE, in a task $e_{k}$ with embedding $c_{k}$, the positive skill $\pi_{c_{k}}^{+}$ conditions on $c_{k}$ and generates positive trajectories $\tau_{c_{k}}^{+}$, and the negative skill $\pi_{c_{k}}^{-}$ generates negative trajectories $\tau_{c_{k}}^{-}$. The top graphs show the relationship between $c$, $\pi_{c}$ and $\tau_{c}$.
4.3 Skill-aware noise contrastive estimation: a tighter $K$-sample estimator
Despite InfoNCE’s success as a $K$-sample estimator for approximating MI [Laskin et al., 2020, Eysenbach et al., 2022], its learning efficiency plunges due to limited numerical precision, which is called the $\log$-$K$ curse, i.e., $I_{\text{InfoNCE}}\leq\log K\leq I_{\text{SaMI}}\leq I(c;\tau_{c})$ [Chen et al., 2021] (see proof in Appendix B). When $K\to+\infty$, we can expect $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ [Guo et al., 2022]. However, increasing $K$ is too expensive, especially in complex environments with enormous negative sample space. In response, we propose a novel $K$-sample estimator with a reduced required sample space size $K^{*}$ ($K^{*}\ll+\infty$). First, we define $K^{*}$:
Definition 2 ($\mathbf{K^{*}}$)
$K^{*}=|c|\cdot|\pi_{c}|\cdot M$ is defined as the number of trajectories in the replay buffer (i.e., the sample space), in which $|c|$ represents the number of different context embeddings $c$, $|\pi_{c}|$ represents the number of different skills $\pi_{c}$, and $M$ is a natural number.
Figure 5: A practical framework for SaNCE. The context encoder embeds the trajectories into context embeddings, in which the queries are generated by context encoder $\psi$ and the negative/positive embeddings are produced by the momentum context encoder $\psi^{*}$. The encoder is updated with gradients from both the SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and the RL loss $\mathcal{L}_{RL}$. The replay buffer is task-specific.
Note that $K^{*}$ is the maximum batch size that can be sampled in contrastive learning. Therefore, to ensure that $I_{\text{InfoNCE}}$ is a tight bound of $I_{\text{SaMI}}$, we require that $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ when $K\to K^{*}$. Under the definition of $K^{*}$, the replay buffer can be divided according to the different context embeddings $c$ and skills $\pi_{c}$ (i.e., observing context embeddings $c$ and skills $\pi_{c}$). In real-world robotic control tasks, the sample space size significantly increases due to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. Taking the sample space of InfoNCE as an example (Figure 4(a)), in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{c_{1}}$ generated after executing the skill $\pi_{c_{1}}$ in task $e_{1}$, and negative samples are trajectories $\{\tau_{c_{2}},...\}$ from other tasks $\{e_{2},...\}$. The permutations and combinations of $N$ environmental features lead to an exponential growth in task number $|c|$, which in turn results in an increase of sample space $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M$.
We introduce a tight $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), which is used to approximate $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ with a reduced $K^{*}_{\text{SaNCE}}$. For SaNCE, both positive samples $\tau_{c_{1}}^{+}$ and negative samples $\tau_{c_{1}}^{-}$ are sampled from the current tasks $e_{1}$, but are generated by executing positive skills $\pi_{c_{1}}^{+}$ and negative skills $\pi_{c_{1}}^{-}$, respectively. Here, a positive skill is intuitively defined by whether it is optimal for the current task $e$, with a more formal definition provided in Section 4.4. For instance, in a cube-moving task under a large friction setting, the agent executes a skill $\pi_{c}^{+}$ after several iterations of learning, and obtains corresponding trajectories $\tau_{c}^{+}$ where the cubes leave the table surface. This indicates that the skill $\pi_{c}^{+}$ is Pick&Place and other skills $\pi_{c}^{-}$ may include Push or Flip (flipping the cube to the goal position), with corresponding trajectories $\tau_{c}^{-}$ where the cube remains stationary or rolls on the table. Formally, we can optimise the $K$-sample lower bound $I_{\text{SaNCE}}$ to approximate $I_{\text{SaMI}}$:
$\quad I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$
$=\mathbb{E}_{p(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K}^{-})}\left[\log\left(\frac{K\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
$\leq I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ (3)
where $f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}})=e^{\psi(\tau_{c_{1}})^{\top}\cdot\psi^{*}(\tau_{c_{1}})/\beta}$. The query $c_{1}=\psi(\tau_{c_{1}})$ is generated by the context encoder $\psi$. For training stability, we use a momentum encoder $\psi^{*}$ to produce the positive and negative embeddings. SaNCE significantly reduces the required sample space size $K^{*}_{\text{SaNCE}}$ by sampling trajectories $\tau_{c}$ based on different skills $\pi_{c}$ (Figure 4(b)) in task $e_{1}$, so that $K^{*}_{\text{SaNCE}}=|c|\cdot|\pi_{c}|\cdot M=|\pi_{c_{1}}|\cdot M\leq K^{*}_{\text{InfoNCE}}$ ($|c|=|c_{1}|=1$). Therefore, $I_{\text{SaNCE}}$ satisfies Lemma 2:
Lemma 2
With a skill-aware context encoder $\psi$, we have $I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq\log K\leq I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$. (see proof in Appendix B)
SaNCE is a plug-and-play module for any other NCE. For example, the negative sample space can be more diverse by combining SaNCE and InfoNCE with $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M$ (Figure 4(c)). A detailed analysis of the sample size of $I_{\text{Sa+InfoNCE}}$ can be found in Appendix C.
4.4 Optimal/Sub-optimal skill-aware sampling strategy
In this section, we further define the positive and negative skills and propose a practical trajectory sampling method. SaNCE requires sample trajectories by ovserving skills $\pi_{c}$. However, we cannot access the prior distribution $p(\pi_{c})$, making it impossible to enumerate all distinct skills $\pi_{c}$. Therefore, in the optimal/sub-optimal skill-aware sampling method, positive skills $\pi_{c}^{+}$ are defined as the optimal skills achieving the highest return $\sum_{i=t}^{t+T}R(s_{i},a_{i})$ in task $e$, while negative skills $\pi_{c}^{-}$ encompass all the other sub-optimal skills with lower returns in task $e$. Thus, we can simply sample the trajectory with the ranked highest return from the replay buffer as the positive sample $\tau_{c}^{+}$, and the one with the lowest return as the negative sample $\tau_{c}^{-}$. The SaNCE loss is then minimised to bring the context embeddings of the highest return trajectories closer in the embedding space while distancing those of negative trajectories.
Note that, at the end of the training, the top-ranked trajectories in the ranked replay buffer correspond to positive samples $\tau_{c}^{+}$ with high returns, and the lower-ranked ones are negative samples $\tau_{c}^{-}$ with low returns. However, before the agent is able to achieve high returns, all trajectories are with low returns. Therefore, our SaNCE loss is a soft version of the $K$-sample SaNCE estimator:
$\mathcal{L}_{\text{SaNCE}}=-\max\left(||\psi(\tau_{c}^{+}),\psi(\tau_{c}^{-})||_{L2},1\right)\cdot I_{\text{SaNCE}}$ (4)
where $||\cdot||_{L2}$ represents the Euclidean distance [Tabak, 2014]. Figure 5 provides a practical framework of SaNCE, with a cube-moving example task $e_{1}$ under high friction. In task $e_{1}$, the positive skill $\pi_{c_{1}}^{+}$ is the Pick&Place skill, which is used to generate queries $\psi(\tau_{c_{1}}^{+})$ and positive embeddings $\psi^{*}(\tau_{c_{1}}^{+})$; after executing Push skill we get negative samples $\tau_{c_{1}}^{-}$ and negative embeddings $\psi^{*}(\tau_{c_{1}}^{-})$.
5 Experiments
We use a three-step process to demonstrate the benefits of SaMI in each environment and answer three questions: (1) Does optimising SaMI lead to increased returns during training and zero-shot generalisation (see Table 5.1 and 2)?; (2) Does SaMI help the RL agents to be versatile and embody multiple skills (see Figure 6)?; (3) Can SaNCE overcome the $\log$-$K$ curse in sample-limited scenarios (see Table 5.1 and 2, and Section 5.4)?
5.1 Experimental setup
Modified benchmarks with multiple environmental features.  [^1] Our modified benchmarks are open-sourced at https://github.com/uoe-agents/Skill-aware-Panda-gym We demonstrate the efficacy of our method using two benchmarks, Panda-gym [Gallouédec et al., 2021] and MuJoCo [Todorov et al., 2012] (details in Sections 5.2 and 5.3). The benchmarks are modified to be influenced by multiple environmental features simultaneously. Environmental features are sampled at the start of each episode during both training and testing phases. During training, we uniform-randomly select a combination of environmental features from a training set. At test time, we evaluate each algorithm in unseen tasks with different environmental features outside the training range. Generalisation performance is measured in two different regimes: moderate and extreme. The moderate regime draws environmental features from a closer range to the training range compared to the extreme. For all our experiments, we report the mean and standard deviation of the models trained over five seeds in both training and test tasks. Further experimental details are available in Appendix D.
Baselines. In our experiments, we primarily compare InfoNCE with SaNCE to demonstrate the performance improvements brought by SaNCE. Therefore, we compare our method with three prevailing and competitive baselines. First, we consider CCM [Fu et al., 2021], which is equipped with InfoNCE. Additionally, we consider TESAC [Yu et al., 2020], which employs a value function loss, allowing us to evaluate the impact on the context encoder without contrastive loss. Given that CCM and TESAC are using RNN encoder, we also consider PEARL [Rakelly et al., 2019], which utilises an MLP context encoder and a similar loss as TESAC.
Our methods.  [^2] Our code and experimental data are available at https://github.com/uoe-agents/SaMI We employ Soft Actor-Critic (SAC) [Haarnoja et al., 2018] as the base RL algorithm. SaNCE can be integrated with any Meta-RL algorithm. Hence, we equip two RL algorithms with SaNCE: (1) SaTESAC is TESAC with SaNCE, which solely uses SaNCE for contrastive learning, with a significantly smaller sample space than that of other algorithms, as shown in Figure 4(b); (2) SaCCM is CCM with SaNCE, where the contrastive learning combines InfoNCE and SaNCE, as shown in Figure 4(c). The purpose is to compare both the impact of different negative sample spaces on the contrastive learning and downstream RL tasks, and the adaptability of SaNCE across different algorithms. In Appendix G, we provide a comparison in MuJoCo with numerical results from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] because of the exact same environmental setting. The sample spaces of PEARL, TESAC, and CCM are all $|c|$ times larger than that of SaTESAC.
5.2 Panda-gym
Task description. Our modified Panda-gym benchmark contains a robot arm control task using the Franka Emika Panda [Gallouédec et al., 2021], where the robot needs to move a cube to a target position. Unlike previous works, the robot can flexibly execute different skills (Push and Pick&Place) to solve tasks. In our experiments, we simultaneously modify multiple environmental features (cube mass and table friction) that characterise the transition dynamics. In some environmental settings,
Table 1: Comparison of success rate with baselines in Panda-gym (over 5 seeds).The blue background indicates that our method surpasses PEARL, TESAC and CCM. Bold text signifies the highest average return.
Training Test (moderate) Test (extreme)
PEARL 0.42$\pm$0.19 0.10$\pm$0.06 0.11$\pm$0.05
TESAC 0.64$\pm$0.16 0.33$\pm$0.23 0.27$\pm$0.17
CCM 0.82$\pm$0.04 0.50$\pm$0.23 0.29$\pm$0.23
SaTESAC 0.92$\pm$0.05 0.55$\pm$0.25 0.40$\pm$0.35
SaCCM 0.94$\pm$0.02 0.58$\pm$0.27 0.46$\pm$0.36
only a specific skill can accomplish the task. For example, an agent needs to learn the Pick&Place skill to avoid high friction. Thus, the agent must determine when and which skill to execute.
Evaluating zero-shot generalisation. As shown in Table 5.1, SaTESAC and SaCCM achieve superior generalisation performance compared to PEARL, TESAC, and CCM, with a smaller sample space. In training, moderate test, and extreme test tasks, SaTESAC achieved up to 43.75%, 66.67%, and 48.15% higher success rates, respectively, compared to TESAC, while SaCCM improved success rates by 14.63%, 16.00%, and 58.62% compared to CCM.
Figure 6: (a) t-SNE visualisation of context embeddings in the Panda-gym environment. The clustered points in the yellow box correspond to the Push skill executed in scenarios with a large mass. Heatmap of (b) success rate, and probability of learned (c) Push skill and (d) Pick&Place skill of SaCCM. Under large mass scenarios, the probability of executing Push skill is higher than Pick&Place skill.
Skill analysis. We visualise the context embedding of SaCCM using t-SNE [Van der Maaten and Hinton, 2008]. We determine the skills by detecting contact points between the end effector and the cube and between the cube and the table (see Appendix D for more details), and then employ Heatmaps [Waskom, 2021] to visualise skills executed by the agent. As shown in Figure 6, we found that when the cube mass is large (30 Kg and 10 Kg), the agent learned the Push skill (corresponding to the clustered points in the yellow bounding box in Figure 6(a). With smaller masses, the agent learned the Pick&Place skill. However, as shown in Figure 15 in Appendix F.1, CCM did not exhibit clear skill grouping. This indicates that SaMI extracts high-quality skill-related information from the trajectories and aids in embodying diverse skills. We provide additional visualisation based on both t-SNE and PCA [Jolliffe, 2002], and heatmap results in Appendix F.
5.3 MuJoCo
Task description. The modified MuJoCo benchmark [Mu et al., 2022, Lee et al., 2020] contains 6 typical robotic control environments based on the MuJoCo physics engine [Todorov et al., 2012]. Hopper, Half-cheetah, Ant, and SlimHumanoid are affected by multiple continuous environmental features (i.e., mass, damping) that characterise the transition dynamics. Both Crippled Ant and Crippled Half-cheetah are more difficult because they are also affected by a discrete environmental feature as one of the actuators in the legs is randomly crippled to change the transition dynamics.
Table 2: Comparison of average return with baselines in modified MuJoCo benchmark (over 5 seeds). The blue background indicates that our method surpasses PEARL, TESAC and CCM. Bold text signifies the highest average return.
Crippled Ant Crippled Half-cheetah
Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
PEARL 182$\pm$73 96$\pm$21 88$\pm$31 2538$\pm$783 1028$\pm$445 908$\pm$185
TESAC 485$\pm$83 129$\pm$32 117$\pm$32 4512$\pm$737 1708$\pm$360 1550$\pm$274
CCM 814$\pm$23 146$\pm$13 158$\pm$31 4153$\pm$569 1417$\pm$710 995$\pm$188
SaTESAC 568$\pm$70 176$\pm$9 196$\pm$17 5380$\pm$917 2334$\pm$169 1848$\pm$341
SaCCM 719$\pm$82 215$\pm$6 219$\pm$9 6210$\pm$402 2656$\pm$231 2070$\pm$651
Ant Half-cheetah
Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
PEARL 153$\pm$63 73$\pm$25 39$\pm$40 1802$\pm$773 530$\pm$270 346$\pm$102
TESAC 761$\pm$234 360$\pm$23 276$\pm$63 5817$\pm$3064 2843$\pm$1303 1074$\pm$446
CCM 1268$\pm$362 1078$\pm$492 548$\pm$177 6839$\pm$542 3053$\pm$347 1422$\pm$332
SaTESAC 911$\pm$54 590$\pm$111 552$\pm$119 7303$\pm$1353 4077$\pm$424 1723$\pm$88
SaCCM 1015$\pm$60 629$\pm$126 588$\pm$92 7662$\pm$316 4384$\pm$223 1828$\pm$52
SlimHumanoid Hopper
Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
PEARL 6947$\pm$3541 3697$\pm$2674 2018$\pm$907 934$\pm$242 874$\pm$366 799$\pm$298
TESAC 9095$\pm$894 7180$\pm$711 3709$\pm$35 1530$\pm$39 1476$\pm$9 1491$\pm$79
CCM 7297$\pm$790 5856$\pm$290 3501$\pm$877 1488$\pm$18 1420$\pm$46 1451$\pm$76
SaTESAC 10501$\pm$2102 8483$\pm$1502 6669$\pm$883 1500$\pm$24 1452$\pm$36 1449$\pm$15
SaCCM 9738$\pm$92 8389$\pm$231 5606$\pm$2027 1472$\pm$53 1456$\pm$13 1446$\pm$94
Results. Table 2 shows the average return of our method and baselines on both training and test tasks. Incorporating SaNCE into TESAC resulted in SaTESAC’s average return increasing by 19.41% in training tasks across all environments, 39.70% in moderate test tasks, and 65.40% in extreme test tasks. Similarly, adding SaNCE to CCM led to SaCCM achieving up to 31.69%, 44.83%, and 40.88% higher average returns in training tasks, moderate test tasks, and extreme test tasks, respectively. Overall, SaTESAC and SaCCM outperform baselines in both training and testing in most of the tasks, in which the Ant, Crippled Ant and Hopper tasks are exceptions. Comparing the visualisation of context embeddings for CCM and SACCM in the Ant, Crippled Ant, and Hopper environments (Appendix F.2), introducing SaNCE does not modify the grouping of context embeddings across different tasks. Notably in Hopper environment, all algorithms maintain consistent average returns when zero-shot generalising to unseen test tasks, with no clear clustering of context embeddings across different tasks (Table 22 in Appendix F.2). This indicates that a single skill may generalise across numerous tasks within the Hopper environment. Hence, we can draw an insight that SaMI is more suitable for environments that require multiple skills to adapt to different environmental features.
5.4 Analysis of the $\log$-$K$ curse in sample-limited scenarios
In this section, we further analyse whether SaNCE can overcome the $\log$-$K$ curse. During the training phases, we sample the environmental features at the beginning of each episode. Therefore, throughout the training process, the context encoder needs to learn the context embedding distribution of multiple tasks. Because InfoNCE requires sampling negative samples across all tasks, and SaNCE samples negative samples from the current task, SaNCE’s negative sample space is $|c|$ times smaller than that of InfoNCE. For example, in the SlimHumanoid environment, where both mass and damping have five values, there can be a maximum of 25 different tasks, making the sampling space of InfoNCE potentially 25 times larger than that of SaNCE. According to Table 5.1 and 2, RL algorithms equipped with SaNCE can achieve better or comparable performance with a much smaller number of negative samples $K$ than InfoNCE. This indicates that SaNCE indeed plays a significant role in addressing the $\log$-$K$ curse, and the SaMI objective helps the contrastive context encoder extract information that is crucial for downstream RL tasks.
Furthermore, the number of negative samples $K$ also relates to two hyperparameters: buffer size, which directly determines the size of the negative sample space; and contrastive batch size, which determines the number of samples used for training contrastive context encoder at each update step. Therefore, we conducted further analysis on these two hyperparameters. As shown in Figure 7, we find that reductions in buffer size and contrastive batch size do not significantly decrease the average return for SaCCM and SaTESAC, which exhibit state-of-the-art performance with small buffers and contrastive batch sizes. Note that the results in Table 2 correspond to a buffer size of 100,000 and a contrastive batch size of 12. Experimental results for all environments (Appendix E.2) further demonstrate that SaNCE is not highly sensitive to changes in $K$, and indeed shows great potential in overcoming the $\log$-$K$ curse.
6 Conclusion and future work
In this paper, we propose a Skill-aware Mutual Information (SaMI) to learn context embeddings for zero-shot generalisation in downstream RL tasks, and a Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI and overcome the $\log$-$K$ curse. RL algorithms equipped with SaMI have achieved state-of-the-art performance in the MuJoCo and Panda-gym benchmarks. Through visual analysis of context embeddings and heatmap analysis of learned skills in Panda-gym, we confirm that SaMI assists the RL agent in embodying diverse skills. These experimental observations indicate that the context encoder, learned by maximising SaMI, helps the RL agent generalise across various tasks by executing different skills. Notably, the optimisation process of SaNCE utilises a far smaller negative sample space than baselines. Coupled with further experimental analysis of buffer size and contrastive batch size on MuJoCo, we demonstrate that SaNCE helps overcome the $\log$-$K$ curse.
Figure 7: Changes of average return in the SlimHumanoid environment. (a) The impact of buffer size on RL control performance (i.e., TESAC, CCM, SaTESAC, SaCCM). (b) The impact of contrastive batch size on the contrastive context encoder (i.e., CCM, SaTESAC, SaCCM).
Our results indicate the importance of aligning the optimisation objectives of representation learning and downstream optimal control to enhance task performance and improve data efficiency.
Given that environmental features are often interdependent, such as a cube’s material correlating with friction and mass, SaMI does not introduce independence assumptions like DOMINO [Mu et al., 2022]. Therefore, future work will focus on verifying and enhancing SaMI’s potential in more complex tasks where environmental features are correlated. This will contribute to our ultimate goal: developing a generalist and versatile agent capable of working across multiple tasks and even real-world tasks in the near future.
Appendix A Proof of Lemma 1
Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, A $K$-sample InfoNCE estimator is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$. InfoNCE is obtained by comparing positive pairs $(x,y_{1})$ and negative pairs $(x,y_{k})$, where $y_{k}\sim y_{2:K}$, as follows:
$I_{\text{InfoNCE}}(x;y|\psi,K)$ $=\mathbb{E}_{p(x,y_{1})p(y_{2:K})}\left[\log\left(\frac{f_{\psi}(x,y_{1})}{\frac{1}{K}\sum_{k=1}^{K}f_{\psi}(x,y_{k})}\right)\right]$ (5)
Step 1. Let us prove the $K$-sample InfoNCE estimator is upper bounded by $\log K$. According to Mu et al. [2022], $\frac{f_{\psi}(x,y_{1})}{\sum_{k=1}^{K}f_{\psi}(x,y_{k})}=\frac{f_{\psi}(x,y_{1})}{f_{\psi}(x,y_{1})+\sum_{k=2}^{K}f_{\psi}(x,y_{k})}\leq 1$. So we have:
$I_{\text{InfoNCE}}(x;y|\psi,K)$ $=\mathbb{E}_{p(x,y_{1})p(y_{2:K})}\left[\log\left(\frac{f_{\psi}(x,y_{1})}{\frac{1}{K}\sum_{k=1}^{K}f_{\psi}(x,y_{k})}\right)\right]$
$=\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p(y_{2:K})}\log\left(\frac{K\cdot f_{\psi}(x,y_{1})}{\sum_{k=1}^{K}f_{\psi}(x,y_{k})}\right)\right]$
$\leq\log K$ (6)
Hence, we have $I_{\text{InfoNCE}}(x;y|\psi,K)\leq\log K$.
Step 2. We have the $I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K)$ which are rightmost and the leftmost side in Lemma 1 according to:
Proposition 1
[Poole et al., 2019] A $K$-sample estimator is an asymptotically tight lower bound to the MI, i.e.,
$I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K),\lim_{x\to+\infty}I_{\text{InfoNCE}}(x;y|\psi,K)\to I(x;y)$
Proof. See Poole et al. [2019] for a neat proof of how the multi-sample estimator (e.g., InfoNCE) lower bounds MI.
Step 3. In this research, the context encoder $\psi$ in $f_{\psi}(x,y)$ is implemented using an RNN to approximate $\frac{p(y|x)}{p(y)}$ [Oord et al., 2019]. For most deep learning platforms, with a powerful learner for $\psi$ and a limited size $K$ such that $I(x;y)\geq\log K$, we can reasonably expect $I_{\text{InfoNCE}}\approx\log K$ after a few training epochs. Therefore, during training, when $K\ll+\infty$, we always have $I(x;y)\geq\log K$.
Proof. See Chen et al. [2021] for more detailed proof.
Step 4. Let us prove that the $K$-sample InfoNCE bound is asymptotically tight. The specific choice of context encoder $\psi$ relates to the tightness of the $K$-sample NCE bound. InfoNCE [Oord et al., 2019] sets $f_{\psi}(x,y)\propto\frac{p(y|x)}{p(y)}$ to model a density ratio which preserves the MI between $x$ and $y$, where $\propto$ stands for ‘proportional to’ (i.e. up to a multiplicative constant). Let us plug in $f_{\psi}(x,y)=f_{\psi}^{*}(x,y)=\frac{p(y|x)}{p(y)}$ into InfoNCE, and we have
$I_{\text{InfoNCE}}(x;y|\psi,K)$ $=\mathbb{E}\left[\log\left(\frac{f_{\psi}^{*}(x,y_{1})}{\sum_{k=1}^{K}f_{\psi}^{*}(x,y_{k})}\right)\right]+\log K$
$=-\mathbb{E}\left[\log\left(1+\frac{p(y)}{p(y|x)}\sum_{k=2}^{K}\frac{p(y_{k}|x)}{p(y_{k})}\right)\right]+\log K$
$\approx-\mathbb{E}\left[\log\left(1+\frac{p(y)}{p(y|x)}(K-1)\mathbb{E}_{y_{k}\sim p(y)}\frac{p(y_{k}|x)}{p(y_{k})}\right)\right]+\log K$
$=-\mathbb{E}\left[\log\left(1+\frac{p(y_{1})}{p(y_{1}|x)}(K-1)\right)\right]+\log K$
$\approx-\mathbb{E}\left[\log\frac{p(y)}{p(y|x)}\right]-\log(K-1)+\log K$
$=I(x;y)-\log(K-1)+\log K$ (7)
Now taking $K\to+\infty$, the last two terms cancel out.
Putting it together. Combining $I(x;y)\geq\log K$ with Proposition 1 and Equation 6, we have Lemma 1:
${I}_{\text{InfoNCE}}(x;y|\psi,K)\leq\log K\leq I(x;y).$ (8)
Besides, according to Equation 7, with sample size $K\to+\infty$, $K$-sample InfoNCE bound is sharp and approaches to the true MI $I(x;y)$, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K\approx I(x;y)$.
Appendix B Proof for Lemma 2
Step 1. According to Lemma 1, we have ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$ (shown in Figure 3).
Step 2. Let us prove SaNCE is a $K$-sample SaNCE estimator and upper bounded by ${\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$. Because $\frac{f_{\psi}(c,\pi_{c},\tau_{c}^{+})}{f_{\psi}(c,\pi_{c},\tau_{c}^{+})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c,k}^{-})}\leq 1$ [Mu et al., 2022], so we have:
$\quad{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
$=\mathbb{E}_{p(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K}^{-})}\left[\log\left(\frac{K\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
$=\mathbb{E}_{p(c_{1},\pi_{c_{1}})}\left[\mathbb{E}_{p(\tau_{c_{1},2:K}^{-})}\log\left(\frac{K\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
$\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$ (9)
Hence, we have ${\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$ (similar to Equation 6).
Step 3. With the definition of $K^{*}$, we can prove ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$. In task $e_{1}$ with context embedding $c_{1}$, SaNCE obtains positive and negative samples from the current task $e_{1}$. Hence the variable $c=c_{1}$ is constant, we have:
$\quad{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
$=\mathbb{E}_{p(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K}^{-})}\left[\log\left(\frac{K\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
$\leq\mathbb{E}_{p(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K^{*}_{\text{SaNCE}}}^{-})}\left[\log\left(\frac{K^{*}_{\text{SaNCE}}\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K^{*}_{\text{SaNCE}}}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
$=\mathbb{E}_{p(\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K^{*}_{\text{SaNCE}}}^{-})}\left[\log\left(\frac{K^{*}_{\text{SaNCE}}\cdot f_{\psi}(\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K^{*}_{\text{SaNCE}}}f_{\psi}(\pi_{c_{1}},\tau_{{c_{1}},k}^{-})}\right)\right]\text{($c_{1}$ is constant.)}$.)
$\approx{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K^{*}_{\text{SaNCE}}}$ (10)
The required sample size $K^{*}_{\text{SaNCE}}=|c_{1}|\cdot|\pi|\cdot M=|\pi|\cdot M$. With $K\to K^{*}_{\text{SaNCE}}$, ${\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$. Correspondingly, for InfoNCE, in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{1}$ generated after executing the skill $\pi_{1}$ in task $e_{1}$, and negative samples are trajectories $\{\tau_{c_{2}}^{-},...\}$ from other tasks $\{e_{2},...\}$. Under the definition of $K^{*}$, we have:
$\quad{\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
$=\mathbb{E}_{p(c,\pi_{c},\tau_{c_{1}})p({\tau_{c_{2:|c|},2:K}})}\left[\log\left(\frac{K\cdot f_{\psi}(c,\pi_{c},\tau_{c_{1}})}{f_{\psi}(c,\pi_{c},\tau_{c_{1}})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c_{2:|c|},k})}\right)\right]$
$\leq\mathbb{E}_{p(c,\pi_{c},\tau_{c_{1}})p(\tau_{c_{2:|c|},2:K^{*}_{\text{InfoNCE}}})}\left[\log\left(\frac{K^{*}_{\text{InfoNCE}}\cdot f_{\psi}(c,\pi_{c},\tau_{c_{1}})}{f_{\psi}(c,\pi_{c},\tau_{c_{1}})+\sum_{k=2}^{K^{*}_{\text{InfoNCE}}}f_{\psi}(c,\pi_{c},\tau_{c_{2:|c|},k})}\right)\right]$
$\approx{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K^{*}_{\text{InfoNCE}}},$ (11)
where $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M\approx|c|\cdot K^{*}_{\text{SaNCE}}$. In real-world robotic control tasks, the sample space size significantly increases due to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. $|c|$ is the number of different tasks and increases exponentially due to the permutations and combinations of $N$ environmental features. When current task $e_{1}$ has context embedding $c_{1}$, the $c_{2:|c|}$ refer to the context embedding for the other tasks. With $K\to K^{*}_{\text{InfoNCE}}$, ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$. Hence, during the training process, ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
Step 4. What remains is to show ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$. The MI for three variables is also called interaction information. According to the definition in McGill [1954], the SaMI can be presented as:
Proposition 2
For the case of three variables, the MI could be written as ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}={\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}-I(c;\tau_{c}|\pi_{c})$.
Using this proposition, one can see that in the case of three variables, interaction information quantifies how much the information shared between two variables differs from what they share if the third variable is known. Several properties of interaction information in the case of three variables have been studied in the literature. Specifically, Yeung [1991] showed that
$-\min\{I(x;\tau_{c}|\pi_{c}),I(c;\pi_{c}|\tau_{c}),I(\tau_{c};\pi_{c}|c)\}\leq I(c;\tau_{c};\pi_{c})\leq\min\{I(c;\tau_{c}),I(c;\pi_{c}),I(\tau_{c};\pi_{c})\}$ (12)
Combining Equation 12 and Proposition 2, we have ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$.
Putting it together. Hence, we have Lemma 2: we always have ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$ (shown in Figure 3), while learning a skill-aware context encoder $\psi$ with SaNCE estimator. $K^{*}_{\text{SaNCE}}<<K^{*}_{\text{InfoNCE}}$ so that $I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$ is a much tighter lower bound of the true $I_{\text{SaMI}(c;\pi_{c};\tau_{c})}$ than ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$.
Appendix C Sample size of $I_{\text{Sa+InfoNCE}}$
We illustrate the sample size of $I_{\text{Sa+InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$ in this section. Sa+InfoNCE plugs SaNCE into InfoNCE, and has positive samples $\tau_{c_{1}}^{+}$ from task $e_{1}$ after executing skill $\pi_{c_{1}}^{+}$, and negative samples are trajectories $\tau_{c_{1:K}}^{-}$ from executing skills $\pi_{c_{1:K}}^{-}$ in task $e_{1:K}$, respectively. Therefore, this is equivalent to observing the variable $c$, then observing the variable $\pi_{c}$, i.e., sampling from the distribution $p(\pi_{c},\tau_{c}|c)p(c)$. We have:
$I_{\text{Sa+InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$
$=\mathbb{E}_{p(c)p(\pi_{c_{1}}^{+},\tau_{c_{1}}^{+}|c)p\left(\left(\pi_{2:|c|}^{-},\tau_{2:|c|}^{-}\right)_{2:K}\right)}\left[\log\left(\frac{K\cdot f_{\psi}(c,\pi_{c_{1}}^{+},\tau_{c_{1}}^{+})}{f_{\psi}(c,\pi_{c_{1}}^{+},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{2:|c|,k}^{-},\tau_{2:|c|,k}^{-})}\right)\right]$
$\leq\mathbb{E}_{p(c)p(\pi_{c_{1}}^{+},\tau_{c_{1}}^{+}|c)p\left(\left(\pi_{2:|c|}^{-},\tau_{2:|c|}^{-}\right)_{2:K^{*}_{\text{Sa+InfoNCE}}}\right)}\left[\log\left(\frac{K^{*}_{\text{Sa+InfoNCE}}\cdot f_{\psi}(c,\pi_{c_{1}}^{+},\tau_{c_{1}}^{+})}{f_{\psi}(c,\pi_{c_{1}}^{+},\tau_{c_{1}}^{+})+\sum_{k=2}^{K^{*}_{\text{Sa+InfoNCE}}}f_{\psi}(c,\pi_{2:|c|,k}^{-},\tau_{2:|c|,k}^{-})}\right)\right]$
$\approx{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K^{*}_{\text{Sa+InfoNCE}}}$ (13)
It should be noted that such a combination will increase the size of the negative sample space, i.e., $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M\geq K^{*}_{\text{SaNCE}}$. Therefore, with the same number $K$ of samples, $I_{\text{Sa+InfoNCE}}$ is less precise and looser than $I_{\text{SaNCE}}$. Hence, a trade-off between sample diversity and the precision of the $K$-sample estimator is required.
Appendix D Environmental setup
Figure 8: (a) Modified Panda-gym benchmarks, (b) the training tasks, the moderate test tasks and the extreme test tasks. The moderate test task setting just has a combinatorial interpolation. The extreme test task setting has unseen ranges of environmental features and is an extrapolation.
D.1 Modified Panda-gym
We modified the original Pick&Place task in Panda-gym [Gallouédec et al., 2021] by setting the $z$ dimension (i.e., the desired height) of the cube’s goal position equal to 00  [^3] If $z$ is not equal to 0, Pick&Place skill is always needed to solve tasks. and maintaining the freedom  [^4] In the original Push task, the grippers are blocked to ensure the agent can only push cubes. However, this hinders the agent from learning Pick&Place skills, leading to failure when it encounters an ”unpushable” scenario. of grippers, so the agent can explore whether it is supposed to push the cube or grasp it. Skills in this benchmark are defined as:
  • Pick&Place skill: This skill specifically refers to the agent trying to use the gripper to grasp the cube, pick it up off the table and place it in the goal position. We determine the Pick&Place skill by detecting no contact points between the table and the cube, two contact points between the robot’s end effector and the cube, and the cube’s height being greater than half its width, meaning it is on the table.
  • Push skill: This skill refers to the agent giving a push force to the cube and then the cube slides to the goal position. We confirm the Push skill by detecting that the cube’s height is equal to half its width.
  • Other skills: Besides the Pick&Place and Push skills, any other different modes of behaviour are classified as other skills. For example, with the Flip skill, the agent applies an initial force to the cube, causing it to roll on the surface or leave the table. In this case, we detect that the cube’s height is greater than half its width, and the number of contact points between the cube and the table is greater than zero.
Some elements in the RL framework are defined as the following:
State space: we use feature vectors which contain cube position (3 dimensions), cube rotation (3 dimensions), cube velocity (3 dimensions), cube angular velocity (3 dimensions), end-effector position (3 dimensions), end-effector velocity (3 dimensions), gripper width (1 dimension), desired goal (3 dimensions) and achieved goal (3 dimensions). Environmental features are not included in the state.
Action space: the action space has 4 dimensions, the first three dimensions are the end-effector’s position changes and the last dimension is the gripper’s width change.
During training, we randomly select a combination of environmental features from a training set sampling combinations from sets: $\text{mass}=1.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0\}$; $\text{mass}\in\{1.0,5.0,10.0\}$ and $\text{friction}=0.1$. At test time, we evaluate each algorithm in all tasks from the moderate test task setting, in which $\text{mass}\in\{5.0,10.0\}$ and $\text{friction}\in\{1.0,5.0,10.0\}$ (shown in Figure 8(b)), and all tasks from extreme test task setting: $\text{mass}=30.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0,30.0\}$; $\text{mass}\in\{1.0,5.0,10.0,30.0\}$ and $\text{friction}=30.0$ (shown in Figure 8(b)).
D.2 Modified MuJoCo
We follow exactly the same environmental settings as DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020]. For Hopper, Half-cheetah, Ant, and SlimHumanoid, we use the environments from the MuJoCo physics engine and use implementation available from Clavera et al. [2019b] and Seo et al. [2020], and scale the mass of every rigid link by scale factor $m$, and scale damping of every joint by scale factor $d$. For Crippled Ant and Crippled Half-cheetah, we use implementation available from Seo et al. [2020] and scale the mass of every rigid link by scale factor $m$, scale damping of every joint by scale factor $d$, and randomly select one leg, and make it crippled to change the dynamic transitions.
Figure 9: Six environments in modified MuJoCo benchmark.
Generalisation performance is measured in two different regimes: moderate and extreme, where the moderate draws environment parameters from a closer range to the training range, compared to the extreme. We have provided our settings for training, extreme and moderate test tasks in Table 3.
Table 3: Environment parameters used for MuJoCo benchmark.
Training Test (Moderate) Test (Extrame) Episode Length
Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 1000
Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 1000
Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 500
SlimHumanoid $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80\}$ $d\in\{0.40,0.50,1.70,1.80\}$ 500
Crippled Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Leg: $\{0,1,2\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Leg: $\{3\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Leg: $\{3\}$ 1000
Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Leg: $\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Leg: $\{4,5\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Leg: $\{1\}$ 1000
D.3 Implementation details
In this section, we provide the implementation details for SaMI. Our codebase is built on top of the publicly released implementation Stable Baselines3 by Raffin et al. [2021] as well as the implementation of InfoNCE by Oord et al. [2019]. A public and open-source implementation of SaMI is available at https://github.com/uoe-agents/SaMI.
Base algorithm. We use SAC [Haarnoja et al., 2018] for downstream evaluation of the learned context embedding. SAC is an off-policy actor-critic method that uses the maximum entropy framework for soft policy iteration. At each iteration, SAC performs soft policy evaluation and improvement steps. We use the same SAC implementation across all baselines and other methods. We trained SAC agents for 1.6 million timesteps on Panda-gym and MuJoCo benchmarks (i.e. Half-cheetah, Ant, Crippled Half-cheetah, Crippled Ant, SlimHumanoid). We evaluated trained agents every $10,000$ timesteps on tasks with fixed random seeds.
Table 4: Hyperparameters used in Panda-gym and MuJoCo benchmarks. Most hyperparameter values are unchanged across tasks except for contrastive batch size and SaNCE loss coefficient.
Hyperparameter Value
Replay buffer size 100,000
Contrastive batch size MuJoCo 12, Panda-gym 256
SaNCE loss coefficient $\alpha$ MuJoCo 1.0, Panda-gym 0.01
Context embedding dimension 6
Hidden state dimension 128
Learning rate (actor, critic and encoder) 1e-3
Training frequency (actor, critic and encoder) 128
Gradient steps 16
Momentum context encoder $\psi^{*}$ soft-update rate 0.05
SAC target soft-update rate critic 0.01, actor 0.05
SAC batch size 256
Discount factor 0.99
Optimizer Adam
Encoder architecture. For our method, the context encoder $\psi$ is modelled as a Long Short-Term Memory (LSTM) that produces a 128-dimensional hidden state vector, subsequently processed through a single-layer feed-forward network to generate a 6-dimensional context embedding. The actor and critic both use the same context encoder to embed trajectories. For contrastive learning, SaNCE utilises a momentum encoder $\psi^{*}$ to generate positive and negative context embeddings [Laskin et al., 2020, He et al., 2020]. Formally, denoting the parameters of $\psi$ as $\theta_{\psi}$ and those of $\psi^{*}$ as $\theta_{\psi^{*}}$, we update $\theta_{\psi^{*}}$ by:
$\theta_{\psi^{*}}\leftarrow m\cdot\theta_{\psi}+(1-m)\cdot\theta_{\psi^{*}}.$ (14)
Here $m\in[0,1)$ is a soft-update rate. Only the parameters $\theta_{\psi}$ are updated by back-propagation. The momentum update in Equation 14 makes $\theta_{\psi^{*}}$ evolve more smoothly by having them slowly track the $\theta_{\psi}$ with $m\ll 1$ (e.g., $m=0.05$ in this research). This means that the target values are constrained to change slowly, greatly improving the stability of learning.
Hyperparameters. A full list of hyperparameters is displayed in Table 4.
Hardware. For each experiment run we use a single NVIDIA Volta V100 GPU with 32GB memory and a single CPU.
Appendix E Additional results
E.1 Balance contrastive and RL updates: loss coefficient $\alpha$
While past work has learned hyperparameters to balance the contrastive loss coefficient $\alpha$ relative to the RL objective [Jaderberg et al., 2016, Bachman et al., 2019], we use both the contrastive and RL objectives together with equal weight $\alpha=1.0$ to be optimal for the MuJoCo benchmark, and $\alpha=0.01$ for the Panda-gym benchmark, we also analyse the effect of loss coefficient $\alpha$ for CCM, SaTESAC and SaCCM in MuJoCo (Figure 11) and Panda-gym (Figure 10) benchmarks.
Figure 10: Loss coefficient $\alpha$ analysis of Panda-gym benchmark in training and test (moderate and extreme) tasks.
Figure 11: Loss coefficient $\alpha$ analysis of MuJoCo benchmark in training tasks.
E.2 Result of $\log$-$K$ curse analysis
E.2.1 Buffer size
Figure 12: Comparison of different buffer sizes in MuJoCo benchmark in training tasks (over 5 seeds). Buffer size = 400000, 100000, 10000, and 1000.
E.2.2 contrastive batch size
Figure 13: Comparison of different contrastive batch sizes in MuJoCo benchmark in training tasks (over 5 seeds). contrastive batch size = 512, 128, 16, and 8.
Appendix F Further skill analysis
F.1 Panda-gym
F.1.1 Visualisation of context embedding
We visualise the context embedding via t-SNE [Van der Maaten and Hinton, 2008] (Figure 15) and PCA [Jolliffe, 2002] (Figure 14). We found that when the mass of the cube is high (30 Kg and 10 Kg), the agent learned the Push skill (the yellow bounding box in Figure 5.1(a)), whereas with lower masses, the agent learned the Pick&Place skill. However, as shown in Figure 15(b), CCM did not display clear skill grouping. This indicates that SaMI extracts high-quality skill-related information from the trajectories and helps with embodying diverse skills.
Figure 14: PCA visualization of context embedding extracted from trajectories collected in Panda-gym environments.
Figure 15: t-SNE visualization of context embedding extracted from trajectories collected in Panda-gym environments.
F.1.2 Heatmap of Panda-gym benchmark
Figure 16: Heatmap of success rate and learned skills of SaCCM. For each grid, the $(i,j)^{th}$ location shows the probability of the skills executed in 100 evaluations with $(\text{mass}=i,\text{friction}=j)$. We determine the skills by detecting contact points between the end effector and the cube and between the cube and the table.
This section provides the heatmap results and further analysis of TESAC, CCM, and SaTESAC on the Panda-gym benchmark. Initially, from the heatmap results of SaTESAC and SaCCM (Figure 6), we observed that with higher cube masses (30 and 10 kg), the agent adopted the Push skill (as indicated by the clustered points in the yellow bounding box in Figure 5.1(a)). At lower masses, the agent adopted the Pick&Place skill.
In contrast to CCM, as seen in Figures 16(e-f), CCM predominantly learned the Pick&Place skill, resulting in a decline in success rates for tasks at $mass=30$, as the agent could not lift the cube off the table using the Pick&Place skill, as depicted in Figure 16(d). The visualisation of the context embedding (Figure 15) did not show clear grouping across different tasks.
Finally, TESAC only mastered the Push skill. The Push skill is relatively simpler to learn than the Pick&Place skill, as it does not require the agent to master manipulating fingers to pick up cubes. Consequently, TESAC’s success rate notably decreased in environments with higher friction.
F.2 MuJoCo
We find that SaMI help the RL agents be versatile and embody multiple skills. For example, Figure 17 displays the t-SNE visualization results of four methods in the Crippled Half-cheetah environment for both test (moderate and extreme) and training tasks. Tasks corresponding to different crippled joints on the same leg are clustered together. Further, we have displayed rendered videos (https://github.com/uoe-agents/SaMI) to better demonstrate the different skills learned.
Figure 17: t-SNE and PCA visualisation of context embedding extracted from trajectories collected in Crippled Half-cheetah environment.
Figure 18: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Crippled Ant environment.
Figure 19: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Half-cheetah environment.
Figure 20: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Ant environment.
Figure 21: t-SNE and PCA visualization of context embedding extracted from trajectories collected in SlimHumanoid environment.
Figure 22: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Hopper environment.
Table 5: Sub-standard comparison of average return with baselines in MuJoCo benchmark (over 5 seeds). The blue background indicates that our method surpasses the baseline, and bold text signifies the highest average return. The numerical results for PPO+DOMINO are copied from Mu et al. [2022]; the numerical results for PPO+CaDM, Vanilla+CaDM, and PE-TS+CaDM are copied from Lee et al. [2020].
Ant Half-cheetah
Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
PPO+DOMINO 227$\pm$86 216$\pm$52 2472$\pm$803 1034$\pm$476
PPO+CaDM 268.6$\pm$77.0 228.8$\pm$48.4 199.2$\pm$52.1 2652.0$\pm$1133.6 1224.2$\pm$630.0 1021.1$\pm$676.6
Vanilla+CaDM 1851.0$\pm$113.7 1315.7$\pm$45.5 821.4$\pm$113.5 3536.5$\pm$641.7 1556.1$\pm$260.6 1264.5$\pm$228.7
PE-TS+CaDM 2848.4$\pm$61.9 2121.0$\pm$60.4 1200.7$\pm$21.8 8264.0$\pm$1374.0 7087.2$\pm$1495.6 4661.8$\pm$783.9
SaTESAC 911$\pm$54 590$\pm$111 552$\pm$119 7303$\pm$1353 4077$\pm$424 1723$\pm$88
SaCCM 1015$\pm$60 629$\pm$126 588$\pm$92 7662$\pm$316 4384$\pm$223 1828$\pm$52
SlimHumanoid Handcapped Half-Cheetah
Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
PPO+DOMINO 7825$\pm$1256 5258$\pm$1039 2503$\pm$658 1326$\pm$491
PPO+CaDM 10455.0$\pm$1004.9 4975.7$\pm$1305.7 3015.1$\pm$1508.3 2356.6$\pm$624.3 1454.0$\pm$462.6 1025.0$\pm$296.2
Vanilla+CaDM 1758.2$\pm$459.1 1228.9$\pm$374.0 1487.9$\pm$339.0 2435.1$\pm$880.4 1375.3$\pm$290.6 966.9$\pm$89.4
PE-TS+CaDM 1371.9$\pm$400.0 903.7$\pm$343.9 814.5$\pm$274.8 3294.9$\pm$733.9 2618.7$\pm$647.1 1294.2$\pm$214.9
SaTESAC 10501$\pm$2102 8483$\pm$1502 6669$\pm$883 5380$\pm$917 2334$\pm$169 1848$\pm$341
SaCCM 9738$\pm$92 8389$\pm$231 5606$\pm$2027 6210$\pm$402 2656$\pm$231 2070$\pm$651
Appendix G A comparison with DOMINO and CaDM
In this section, we give a brief comparison between our methods and methods from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] in the MuJoCo benchmark because we are using the exact same environmental setting.
DOMINO [Mu et al., 2022] is based on the InfoNCE $K$-sample estimator. Their implementation, PPO+DOMINO, is a model-free RL algorithm with a pre-trained context encoder. This encoder reduces the demand for large contrastive batch sizes during training by decoupling representation learning for each modality, simplifying tasks while leveraging shared information. However, a pre-trained encoder necessitates a large sample volume, with DOMINO training PPO agents for 5 million timesteps on the MuJoCo benchmark. In contrast, SaTESAC and SaCCM, trained for 1.6 million timesteps without pre-trained encoders, achieve considerably higher average returns across four environments (Table 5). Therefore, it is crucial to focus on extracting mutual information in contrastive learning that directly optimises downstream tasks, integrating rather than segregating representation learning from task performance.
CaDM [Lee et al., 2020] proposes a context-aware dynamics model adaptable to changes in dynamics. Specifically, they utilise contrastive learning to learn context embeddings, and then predict the next state conditioned on them. We copy the numerical results of PPO+CaDM, Vanilla+CaDM, and PE-TS+CaDM from CaDM [Lee et al., 2020] as their environmental setting is identical to ours, where PPO+CaDM is a model-free RL algorithm, while Vanilla+CaDM and PE-TS+CaDM are model-based. The model-free RL approach, PPO+CaDM, is trained for 5 million timesteps on the MuJoCo benchmark. As shown in Table 5, SaTESAC and SaCCM significantly outperform PPO+CaDM. The model-based RL algorithms, Vanilla+CaDM and PE-TS+CaDM, require 2 million timesteps for learning in model-based setups, compared to our fewer samples (i.e., million timesteps). In the Ant environment, Vanilla+CaDM and PE-TS+CaDM achieve higher returns than SaTESAC and SaCCM; similarly, in the Half-cheetah environment, PE-TS+CaDM outperforms them. Results in the SlimHumanoid and Crippled Half-cheetah environments show that skill-aware context embeddings are notably effective. An insight here is that our method outperforms the model-free CaCM approach, but not the model-based one. Therefore, we consider that a model-based approach to SaMI could be an interesting extension for future work.

### v3

@@
 Abstract
-Meta-Reinforcement Learning (Meta-RL) agents can struggle to operate across tasks with varying environmental features that require different optimal skills (i.e., different modes of behaviours).
+Meta-Reinforcement Learning (Meta-RL) agents can struggle to operate across tasks with varying environmental features that require different optimal skills (i.e., different modes of behaviour).
 Using context encoders based on contrastive learning to enhance the generalisability of Meta-RL agents is now widely studied but faces challenges such as the requirement for a large sample size, also referred to as the $\log$-$K$ curse.
@@
 We empirically find that RL agents that learn by maximising SaMI achieve substantially improved zero-shot generalisation to unseen tasks.
-Additionally, the context encoder equipped with SaNCE demonstrates greater robustness to reductions in the number of available samples, thus possessing the potential to overcome the $\log$-$K$ curse.
+Additionally, the context encoder trained with SaNCE demonstrates greater robustness to a reduction in the number of available samples, thus possessing the potential to overcome the $\log$-$K$ curse.
 1 Introduction
-Meta-Reinforcement Learning (Meta-RL) agents often learn policies that do not generalise across tasks in which the environmental features and optimal skills are different [des Combes et al., 2018, Garcin et al., 2024].
-Consider, for example, a set of cube-moving tasks, in which an agent is required to move a cube to a goal position on a table (Figure 1).
-These tasks become challenging if the underlying environmental features vary between different tasks, such as table friction.
-Suppose an agent
+Reinforcement Learning (RL) agents often learn policies that do not generalise across tasks in which the environmental features and optimal skills are different [des Combes et al., 2018, Garcin et al., 2024].
+Consider a set of cube-moving tasks where an agent is required to move a cube to a goal position on a table (Figure 1).
+These tasks become challenging if environmental features, such as table friction, vary between tasks.
+When facing an unknown environment, the agent
 Figure 1: (a) In a cube-moving environment, tasks are defined according to different environmental features.
 (b) Different tasks have different transition dynamics caused by underlying environmental features, hence optimal skills are different across tasks.
-attempts to push a cube across a table covered by a tablecloth and receives environmental feedback showing that the cube has become "unpushable".
-In this scenario, a generalist agent should infer that the table friction is relatively high.
-Consequently, the agent ought to adapt its strategy, opting to lift the cube off the table to avoid friction, rather than continuing to push.
-Recent advances in Meta-RL [Lee et al., 2020, Agarwal et al., 2021, Mu et al., 2022, Dunion et al., 2023b, a] have shown promising potential to perceive and understand environmental features by inferring context embeddings from a small number of interactions in the environment.
-The context embedding is expected to capture the distribution of tasks and efficiently infer new tasks.
-Meta-RL methods then train a policy conditioned on the context embedding to generalise to multiple tasks.
-The context encoder is the key component for capturing the context embedding from recent experiences [Clavera et al., 2019a, Lee et al., 2020], which affects generalisation performance significantly.
-Some Meta-RL algorithms [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022] are equipped with context encoders based on contrastive learning, which uses the InfoNCE lower bound [Oord et al., 2019] to optimise the mutual information (MI) between context embeddings.
-InfoNCE is a $K$-sample MI estimator and is utilised to learn distinct context embeddings for each task.
-Despite significant advancements, integrating contrastive learning with Meta-RL poses several unresolved challenges, of which two are particularly relevant to this research: (i) existing context encoders based on contrastive learning do not distinguish tasks that require different skills; many prior algorithms only pull embeddings of the same tasks together and push those of different tasks apart.
+needs to explore effectively, understand the environment, and adjust its behaviour accordingly within an episode.
+For instance, if the agent tries to push a cube across a table covered by a tablecloth and finds it “unpushable,” it should infer that the table friction is relatively high and adapt by lifting the cube to avoid friction, rather than continuing to push.
+Recent advances in Meta-Reinforcement Learning (Meta-RL) [Lee et al., 2020, Agarwal et al., 2021, Mu et al., 2022, Dunion et al., 2023b, a, McInroe et al., 2024] enable agents to understand environmental features by inferring context embeddings from a small amount of exploration, and to train a policy conditioned on the context embedding to generalise to novel tasks.
+In recent years, unsupervised contrastive learning algorithms have been shown to learn context embeddings that perform remarkably well on generalisation tasks [Clavera et al., 2019a, Lee et al., 2020].
+In particular, some Meta-RL algorithms [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022] train context encoders by maximising InfoNCE [Oord et al., 2019], which has yielded vital insights into contrastive learning through the lens of mutual information (MI) analysis.
+The InfoNCE objective can be interpreted as a $K$-sample lower bound on the MI between trajectories and context embeddings.
+Despite significant advances, integrating contrastive learning with Meta-RL poses several unresolved challenges, of which two are particularly relevant to this research: (i) Existing context encoders based on contrastive learning do not distinguish tasks that require different skills.
+Many prior algorithms only pull embeddings of the same tasks together and push those of different tasks apart.
 However, for example, a series of cube-moving tasks with high friction may only require a Pick&Place skill (picking the cube off the table and placing it at the goal position), making further differentiation unnecessary.
-(ii) K-sample MI estimators are sensitive to the sample size K (i.e., the log-K curse) [Poole et al., 2019]; a substantial quantity of negative samples is required to approximate the true MI, which is challenging due to RL’s low sample efficiency [Franke et al., 2021] and often impractical for achieving accurate MI estimation [Arora et al., 2019, Nozawa and Sato, 2021].
-The effectiveness of $K$-sample MI estimators breaks down with a small sample size and leads to a significant performance drop in downstream RL tasks [Mnih and Teh, 2012, Guo et al., 2022].
-To enhance RL generalisation across different tasks, we propose that the context embeddings should optimise downstream tasks and indicate whether the current skill remains optimal or requires further exploration, thereby addressing issue (i).
-This approach also reduces the necessary sample size and helps to overcome issue (ii) by concentrating solely on extracting task-relevant MI.
-Specifically, we propose a three-step process tailored to RL: (1) We introduce Skill-aware Mutual Information (SaMI), a smaller ground-truth MI that discriminates context embeddings according to skills by maximising the MI between context embedding, skills and trajectories.
-Additionally, we provide a theoretical proof of why introducing skills can make the ground-truth MI smaller, thus making it easier to optimise; (2) We propose a more data-efficient $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), used to optimise SaMI and reduce the negative sample space based on skills to overcome the $log$-$K$ curse; (3) We provide a framework to equip any Meta-RL algorithm with SaNCE in practice, and propose a practical skill-aware sampling method for SaNCE when the skills are not known in advance.
-We demonstrate empirically in MuJoCo [Todorov et al., 2012] and Panda-gym [Gallouédec et al., 2021] that SaMI improves the zero-shot generalisation performance in sets of previously unseen tasks with moderate and extreme difficulty.
-In the MuJoCo benchmark, SaMI enhances two Meta-RL algorithms [Yu et al., 2020, Fu et al., 2021] by achieving an average increase of 42.27% in returns for moderate testing tasks and a 53.14% improvement for extreme testing tasks.
-Meanwhile, in the Panda-gym benchmark, SaNCE boosts success rates by an average of 20.23% in moderate tests and 18.36% in extreme tests.
-This indicates SaMI’s advantage in encoding information that enables an agent to execute effective skills in downstream control tasks.
-SaNCE-based RL algorithms utilise smaller sample spaces while achieving improved downstream control performance, indicating their potential to overcome the $\log$-$K$ curse.
-Visualisation of the learned context embeddings shows distinct clusters corresponding to different skills, indicating that SaMI is a step towards more robust and versatile RL agents.
-Figure 2: A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$.
-Executing different skills (e.g., Push skill $\pi(\cdot|c_{1})$ and Pick&Place skill $\pi(\cdot|c_{2})$) shows different modes of behaviours in trajectories (e.g., $\tau^{+}_{c_{1}}$ where the cube moves on the table; $\tau^{+}_{c_{2}}$ where the cube leaves the table).
+(ii) Existing K-sample MI estimators are sensitive to the sample size K (i.e., the log-K curse) [Poole et al., 2019].
+The low sample efficiency of RL [Franke et al., 2021] and the sample limitations in zero-shot generalisation make collecting a substantial quantity of samples often impractical [Arora et al., 2019, Nozawa and Sato, 2021].
+The effectiveness of $K$-sample MI estimators breaks down with a finite sample size and leads to a significant performance drop in downstream RL tasks [Mnih and Teh, 2012, Guo et al., 2022].
+To enhance RL generalisation across different tasks, we propose that the context embeddings should optimise downstream tasks and indicate whether the current skill remains optimal or requires further exploration.
+This also reduces the necessary sample size by focusing solely on skill-related information.
+In this work, (1) we introduce Skill-aware Mutual Information (SaMI), a generalised form of MI objective between context embeddings, skills, and trajectories, designed to address issue (i).
+We provide a theoretical proof showing that by introducing skills as a third variable into the MI of context embeddings and trajectories, the resulting SaMI is smaller and easier to optimise.
+Furthermore, (2) we propose a data-efficient $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI, effectively addressing issue (ii).
+Additionally, (3) we propose a practical skill-aware trajectory sampling strategy that shows how to sample positive and negative examples without relying on any prior skill distribution.
+In that way, Meta-RL agents autonomously acquire a set of skills applicable to many tasks, with these skills emerging solely from the SaMI learning objective and data.
+We demonstrate empirically in MuJoCo [Todorov et al., 2012] and Panda-gym [Gallouédec et al., 2021] that SaMI enhances the zero-shot generalisation capabilities of two Meta-RL algorithms [Yu et al., 2020, Fu et al., 2021] by achieving higher returns and success rates on previously unseen tasks, ranging from moderate to extreme difficulty.
+Visualisation of the learned context embeddings reveals distinct clusters corresponding to different skills, suggesting that the SaMI learning objective enables the context encoder to capture skill-related information from trajectories and incentivise Meta-RL agents to acquire a diverse set of skills.
+Moreover, SaNCE enables Meta-RL algorithms to use smaller sample spaces while achieving improved downstream control performance, indicating their potential to overcome the $\log$-$K$ curse.
+Figure 2: A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$ (shortened as $\pi_{c}$).
+The policy $\pi$ conditioned on a fixed $c$ alters the state of the environment in a consistent way, thereby exhibiting a mode of skill.
+The skill $\pi(\cdot|c_{1})$ moves the cube on the table in trajectory $\tau^{+}_{c_{1}}$ and is referred to as the Push skill; correspondingly, the Pick&Place skill $\pi(\cdot|c_{2})$ takes the cube off the table and places it in the goal position in the trajectory $\tau^{+}_{c_{2}}$.
 2 Related works
 Meta-RL.
-Meta-RL methods train an agent conditioned on context embeddings to improve generalisation to unseen tasks.
-As the key component of Meta-RL, the quality of the context embedding can significantly affect the agent’s performance.
+By conditioning on an effective context embedding, Meta-RL policies can zero-shot generalise to new tasks with a small amount of exploration Kirk et al.
+[2023].
 Existing algorithms can be categorised into three types based on different context embeddings.
@@
 PEARL [Rakelly et al., 2019] learns probabilistic context embeddings by recovering the value function.
-Multi-task SAC + TE (TESAC) [Yu et al., 2020] uses the Task Embedding (TE) [Hausman et al., 2018] as input to policies, parameterising the learned policies via a shared embedding space and maximising average returns.
+Multi-task SAC with task embeddings as inputs to policies (TESAC) [Hausman et al., 2018, Yu et al., 2020] parameterises the learned policies through a shared embedding space, aiming to maximise the average returns across tasks.
 However, the update signals from the RL loss are stochastic and weak, and may not capture the similarity relations among tasks [Fu et al., 2021].
@@
 Contrastive learning.
-Contrastive learning has been applied to RL due to its significant momentum in representation learning in recent years, attributed to its superior effectiveness [Tishby and Zaslavsky, 2015, Hjelm et al., 2019, Dunion and Albrecht, 2024], ease of implementation [Oord et al., 2019], and strong theoretical connection to mutual information (MI) estimation [Poole et al., 2019].
-MI is often estimated through noise-contrastive estimation (NCE) [Gutmann and Hyvärinen, 2010], InfoNCE [Oord et al., 2019], and variational objectives [Hjelm et al., 2019].
-InfoNCE has garnered recent interest due to its lower variance [Song and Ermon, 2020] and superior performance in downstream tasks.
-However, InfoNCE may underestimate true MI because it is constrained by the number of samples $K$.
-To address this issue, CCM [Fu et al., 2021] leverages InfoNCE to learn the context embedding with a large number of samples.
-DOMINO [Mu et al., 2022] reduces the total MI by introducing an independence assumption; however, this results in biased information.
-Correspondingly, we focus on proposing an unbiased alternative MI objective and a more data-efficient $K$-sample estimator tailored for downstream RL tasks, which, to our knowledge, have not been addressed in previous research.
+Contrastive learning has been applied to RL due to its significant momentum in representation learning in recent years, attributed to its superior effectiveness [Tishby and Zaslavsky, 2015, Hjelm et al., 2019, Dunion and Albrecht, 2024], ease of implementation [Oord et al., 2019], and strong theoretical connection to MI estimation [Poole et al., 2019].
+MI is often estimated using InfoNCE [Oord et al., 2019] that has gained recent attention due to its lower variance [Song and Ermon, 2020] and superior performance in downstream tasks.
+However, InfoNCE may underestimate the true MI when the sample size $K$ is finite.
+To address this limitation, CCM [Fu et al., 2021] uses a large number of samples for maximising InfoNCE.
+DOMINO [Mu et al., 2022] reduces the true MI by introducing an independence assumption; however, this results in biased estimates.
+We focus on proposing an unbiased alternative MI objective and a more data-efficient $K$-sample estimator tailored for downstream RL tasks, which, to our knowledge, have not been addressed in previous research.
 3 Preliminaries
@@
 The Meta-RL agent’s goal is to learn a generalisable policy $\pi$ that is robust to such dynamic changes.
-Specifically, given a set of training tasks $e$ sampled from $\xi_{\text{train}}(e)$, we aim to learn a policy that can maximise the discounted returns, $\arg\max_{\pi}\mathbb{E}_{e\sim\xi_{\text{train}}(e)}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim P_{e}(s_{t+1}|s_{t},a_{t})]$, and can produce accurate control for unseen test tasks $e$ sampled from $\xi_{\text{test}}(e)$.
+Specifically, given a set of training tasks $e$ sampled from $\xi_{\text{train}}(e)$, we aim to learn a policy that can maximise the discounted returns, $\arg\max_{\pi}\mathbb{E}_{e\sim\xi_{\text{train}}(e)}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim P_{e}(s_{t+1}|s_{t},a_{t})]$, and can produce accurate control for unseen test tasks sampled from $\xi_{\text{test}}(e)$.
 Contrastive learning.
-A Meta-RL agent first consumes a collected trajectory $\tau_{t:t+T}=\{s_{t},a_{t},r_{t},s_{t+1},...,s_{t+T},a_{t+T},r_{t+T}\}$ and outputs a context embedding $c$ through a context encoder $\psi(c|\tau_{t:t+T})$, then executes policy $\pi$ conditioned on the current state $s_{t}$ and context embedding $c$ to work across multiple tasks.
-As a key component, a context encoder $\psi$ captures the unknown context $c$ of the task.
-We focus on a context encoder that optimises the InfoNCE bound $I_{\text{InfoNCE}}(x;y)$, which is a $K$-sample estimator of the MI $I(x;y)$ [Oord et al., 2019].
+In Meta-RL, the context encoder $\psi(c|\tau_{c,0:t})$ first takes the trajectory $\tau_{c,0:t}=\{s_{0},a_{0},r_{0},...,s_{t}\}$ from the current episode as input and compresses it into a context embedding $c$ [Fu et al., 2021].
+Then, the policy $\pi$, conditioned on context embedding $c$, consumes the current state $s_{t}$ and outputs the action $a_{t}$.
+As a key component, the quality of context embedding $c$ can affect algorithms’ performance significantly.
+MI is an effective measure of embedding quality [Goldfeld et al., 2019], hence we focus on a context encoder that optimises the InfoNCE objective $I_{\text{InfoNCE}}(x;y)$, which is a $K$-sample estimator and lower bound of the MI $I(x;y)$ [Oord et al., 2019].
 Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, $I_{\text{InfoNCE}}(x;y)$ is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$:
 $I_{\text{InfoNCE}}(x;y|\psi,K)=\mathbb{E}\left[\log\frac{f_{\psi}(x,y_{1})}{\frac{1}{K}\sum_{k=1}^{K}f_{\psi}(x,y_{k})}\right].$ (1)
-InfoNCE constructs a formal lower bound to the MI, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\leq I(x;y)$ [Guo et al., 2022, Chen et al., 2021].
-Given two inputs $x$ and $y$, their embedding similarity is $f_{\psi}(x,y)=e^{\psi(x)^{\top}\cdot\psi(y_{1})/\beta}$, where $\psi$ is the context encoder that projects $x$ and $y$ into the context embedding space, the dot product is used to calculate the similarity score between $\psi(x),\psi(y)$ pairs [Wu et al., 2018, He et al., 2020], and $\beta$ is a temperature hyperparameter that controls the sensitivity of the product.
+InfoNCE constructs a formal lower bound on the MI, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\leq I(x;y)$ [Guo et al., 2022, Chen et al., 2021].
+Given two inputs $x$ and $y$, their embedding similarity is $f_{\psi}(x,y)=e^{\psi(x)^{\top}\cdot\psi(y)/\beta}$, where $\psi$ is the context encoder that projects $x$ and $y$ into the context embedding space, the dot product is used to calculate the similarity score between $\psi(x),\psi(y)$ pairs [Wu et al., 2018, He et al., 2020], and $\beta$ is a temperature hyperparameter that controls the sensitivity of the product.
 Some previous Meta-RL methods [Lee et al., 2020, Mu et al., 2022] learn a context embedding $c$ by maximising $I_{\text{InfoNCE}}(c;\tau_{c}|\psi,K)$ between the context $c$ embedded from a trajectory in the current task, and the historical trajectories $\tau_{c}$ under the same environmental features setting.
 4 Skill-aware mutual information optimisation for Meta-RL
-4.1 The $log$-$K$ curse of $K$-sample MI estimators
-In this section, we provide a theoretical analysis of the challenge inherent in learning a $K$-sample estimator for MI, commonly referred to as the $log$-$K$ curse.
+4.1 The $\boldsymbol{\log}$-$\boldsymbol{K}$ curse of $\boldsymbol{K}$-sample MI estimators
+In this section, we provide a theoretical analysis of the challenge inherent in learning a $K$-sample estimator for MI, commonly referred to as the $\log$-$K$ curse.
 Based on this theoretical analysis, we give insights to overcome this challenge.
-If we have sufficient training epochs for the context encoder $\psi$ in Equation 1 and a sample size $K$, $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K$ [Mnih and Teh, 2012, Guo et al., 2022].
+Given that we focus on the generalisation of RL, we only consider cases with a finite sample size of $K$.
+If a context encoder $\psi$ in Equation (1) has sufficient training epochs, then $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K$ [Mnih and Teh, 2012, Guo et al., 2022].
 Hence, the MI we can optimise is bottlenecked by the number of available samples, formally expressed as:
@@
 […]
-To ensure that the $K$-sample estimator is a tight bound of the ground-truth MI, inspired by Lemma 1, we derive three key insights when learning a context encoder: (1) focus on a ground-truth MI that is smaller than $I(x;y)$ and achievable with a smaller sample size; (2) develop a $K$-sample estimator tighter than $I_{\text{InfoNCE}}$; (3) increasing sample quantity $K$, however, this is usually impractical.
-A meta-RL agent learns a context encoder by maximising MI between trajectories $\tau_{c}$ and context embeddings $c$.
-Driven by insight (1), we introduce Skill-aware Mutual Information (SaMI) in Section 4.2, designed to enhance the zero-shot generalisation of downstream RL tasks.
-Corresponding to insight (2), we propose Skill-aware Noise Contrastive Estimation (SaNCE) to maximise SaMI with limited samples in Section 4.3.
-Finally, Section 4.4 demonstrates how to equip any Meta-RL agent with SaNCE in practice.
-Figure 3: A $K$-sample estimator $I_{\text{InfoNCE}(c;\pi_{c};\tau_{c})}$, with limited sample size $K$, is a loose lower bound of $I(c;\tau_{c})$ and leads to lower performance embeddings.
-$I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ is a lower ground-truth MI, and $I_{\text{SaNCE}}(c;\pi_{c};\tau_{c})$ is a tighter lower bound of $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$.
 4.2 Skill-aware mutual information: a smaller ground-truth MI
-An useful tool in learning a versatile agent is to understand when to explore novel skills or switch between existing skills in multi-task settings. To start with, we define skills:
+We aim for our MI learning objective to incentivise agents to acquire a diverse set of skills, enabling them to generalise effectively across tasks.
+To start with, we define skills [Eysenbach et al., 2018]:
 Definition 1 (Skills)
@@
 If a skill $\pi_{c}$ is conditioned on a state $s_{t}$, we can sample actions $a_{t}\sim\pi(\cdot|c,s_{t})$.
-If we sample actions from $\pi_{c}$ at consecutive timesteps, we gain a trajectory $\tau_{c}=\{s_{t},a_{t},r_{t},s_{t+1},...,s_{t+T},a_{t+T},r_{t+T}\}$.
-After executing skills, an agent should distinguish tasks through environmental feedback to be aware of underlying environmental features $e=\{e^{0},e^{1},...,e^{N}\}$.
-In this regard, the context encoder $\psi$ should learn by maximising the MI between context embedding $c$, skills $\pi_{c}$, and environmental feedback (i.e., trajectories $\tau_{c}$).
-We achieve this learning process by maximising the MI $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$, in which we introduce a variable, skill $\pi_{c}$, into $I(c;\tau_{c})$.
-Formally, we propose a novel MI optimisation objective for a context encoder, Skill-aware Mutual Information (SaMI), which is defined as:
-$I_{\text{SaMI}}(c;\pi_{c};\tau_{c})=\mathbb{E}_{p(c,\pi_{c},\tau_{c})}\left\{\log\ \frac{p(\tau_{c}|c,\pi_{c})}{p(\tau_{c})}\right\}.$ (2)
+After sampling actions from $\pi_{c}$ at consecutive timesteps, we obtain a trajectory $\tau_{c,t:t_{T}}=\{s_{t},a_{t},r_{t},s_{t+1},\ldots,s_{t+T},a_{t+T},r_{t+T}\}$ which demonstrates a consistent mode of behaviour.
+After a limited amount of exploration, an agent should be able to infer the task (i.e., environmental features $e={e^{0},e^{1},\ldots,e^{N}}$) and adapt accordingly within the current episode.
+The context embedding should encompass skill-related information, guiding the policy on when to explore new skills or switch between existing ones.
+We propose that the context encoder $\psi$ should be trained by maximising the MI between the context embedding $c$, skills $\pi_{c}$, and trajectories $\tau_{c}$.
+To this end, we propose a novel MI optimisation objective, Skill-aware Mutual Information (SaMI), defined as:
+$I_{\text{SaMI}}(c;\pi_{c};\tau_{c})=I(c;\tau_{c})-I(c;\tau_{c}|\pi_{c}).$ (2)
+SaMI is defined according to interaction information [McGill, 1954], serving as a generalisation of MI for three variables $\{c,\pi_{c},\tau_{c}\}$.
 Although we cannot evaluate $p(c,\pi_{c},\tau_{c})$ directly, we approximate it by Monte-Carlo sampling, using $K$ samples from $p(c,\pi_{c},\tau_{c})$.
-After introducing the variable $\pi_{c}$, $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ contains less information, i.e., $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$ (see proof in Appendix B).
-So the context encoder $\psi$ has a lower ground-truth MI $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ to approach.
-$I_{\text{SaMI}}$ also enables the RL agent to embody diverse skills and switch skills based on the context embedding $c$.
-In other words, the skill $\pi_{c}$ conditioned on different context embeddings $c$ will exhibit various modes of behaviour, such as the Push skill $\pi_{c_{1}}$ (pushing a cube on the table to the target position) or the Pick&Place skill $\pi_{c_{2}}$.
+As illustrated in Figure 3, a context encoder $\psi$ trained with the objective of maximising $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ converges more quickly, as $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$ (see proof in Appendix B).
+By focusing more on skill-related information, $I_{SaMI}$ enables agents to autonomously discover a diverse range of skills for handling multiple tasks.
 Figure 4: A comparison of sample spaces for task $e_{1}$.
@@
 The top graphs show the relationship between $c$, $\pi_{c}$ and $\tau_{c}$.
-4.3 Skill-aware noise contrastive estimation: a tighter $K$-sample estimator
-Despite InfoNCE’s success as a $K$-sample estimator for approximating MI [Laskin et al., 2020, Eysenbach et al., 2022], its learning efficiency plunges due to limited numerical precision, which is called the $\log$-$K$ curse, i.e., $I_{\text{InfoNCE}}\leq\log K\leq I_{\text{SaMI}}\leq I(c;\tau_{c})$ [Chen et al., 2021] (see proof in Appendix B).
+4.3 Skill-aware noise contrastive estimation: a tighter $\boldsymbol{K}$-sample estimator
+Despite InfoNCE’s success as a $K$-sample estimator for approximating MI [Laskin et al., 2020, Eysenbach et al., 2022], its learning efficiency plunges due to limited numerical precision, which is called the $\log$-$K$ curse, i.e., $I_{\text{InfoNCE}}\leq\log K\leq I_{\text{SaMI}}$ [Chen et al., 2021] (see proof in Appendix B).
 When $K\to+\infty$, we can expect $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ [Guo et al., 2022].
 However, increasing $K$ is too expensive, especially in complex environments with enormous negative sample space.
-In response, we propose a novel $K$-sample estimator with a reduced required sample space size $K^{*}$ ($K^{*}\ll+\infty$).
+To address this, we propose a novel $K$-sample estimator that requires a significantly smaller sample size $K\ll+\infty$.
 First, we define $K^{*}$:
-Definition 2 ($\mathbf{K^{*}}$)
+Definition 2 ($\boldsymbol{K^{*}}$)
 $K^{*}=|c|\cdot|\pi_{c}|\cdot M$ is defined as the number of trajectories in the replay buffer (i.e., the sample space), in which $|c|$ represents the number of different context embeddings $c$, $|\pi_{c}|$ represents the number of different skills $\pi_{c}$, and $M$ is a natural number.
-Figure 5: A practical framework for SaNCE.
-The context encoder embeds the trajectories into context embeddings, in which the queries are generated by context encoder $\psi$ and the negative/positive embeddings are produced by the momentum context encoder $\psi^{*}$.
-The encoder is updated with gradients from both the SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and the RL loss $\mathcal{L}_{RL}$.
-The replay buffer is task-specific.
-Note that $K^{*}$ is the maximum batch size that can be sampled in contrastive learning.
-Therefore, to ensure that $I_{\text{InfoNCE}}$ is a tight bound of $I_{\text{SaMI}}$, we require that $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ when $K\to K^{*}$.
+Figure 5: A practical framework for using SaNCE in the meta-training phase.
+During meta-training, we sample trajectories from the replay buffer for off-policy training.
+Queries are generated by a context encoder $\psi$, which is updated with gradients from both the SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and the RL loss $\mathcal{L}_{RL}$.
+Negative/Positive embeddings are encoded by a momentum context encoder $\psi^{*}$, which is driven by a momentum update with the encoder $\psi$.
+During meta-testing, the meta-trained context encoder $\psi$ embeds the current trajectory, and the RL policy takes the embedding as input together with the state for adaptation within an episode.
+To ensure that $I_{\text{InfoNCE}}$ is a tight bound of $I_{\text{SaMI}}$, we require that $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ when $K\to K^{*}$.
 Under the definition of $K^{*}$, the replay buffer can be divided according to the different context embeddings $c$ and skills $\pi_{c}$ (i.e., observing context embeddings $c$ and skills $\pi_{c}$).
@@
 The permutations and combinations of $N$ environmental features lead to an exponential growth in task number $|c|$, which in turn results in an increase of sample space $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M$.
-We introduce a tight $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), which is used to approximate $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ with a reduced $K^{*}_{\text{SaNCE}}$.
+We introduce a tight $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), which is used to approximate $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ with $K^{*}_{\text{SaNCE}}<K^{*}_{\text{InfoNCE}}$.
 For SaNCE, both positive samples $\tau_{c_{1}}^{+}$ and negative samples $\tau_{c_{1}}^{-}$ are sampled from the current tasks $e_{1}$, but are generated by executing positive skills $\pi_{c_{1}}^{+}$ and negative skills $\pi_{c_{1}}^{-}$, respectively.
@@
 Lemma 2
-With a skill-aware context encoder $\psi$, we have $I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq\log K\leq I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$.
+With a context encoder $\psi$ and finite sample size $K$, we have $I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq\log K\leq I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$.
 (see proof in Appendix B)
-SaNCE is a plug-and-play module for any other NCE.
-For example, the negative sample space can be more diverse by combining SaNCE and InfoNCE with $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M$ (Figure 4(c)).
-A detailed analysis of the sample size of $I_{\text{Sa+InfoNCE}}$ can be found in Appendix C.
-4.4 Optimal/Sub-optimal skill-aware sampling strategy
-In this section, we further define the positive and negative skills and propose a practical trajectory sampling method.
-SaNCE requires sample trajectories by ovserving skills $\pi_{c}$.
-However, we cannot access the prior distribution $p(\pi_{c})$, making it impossible to enumerate all distinct skills $\pi_{c}$.
-Therefore, in the optimal/sub-optimal skill-aware sampling method, positive skills $\pi_{c}^{+}$ are defined as the optimal skills achieving the highest return $\sum_{i=t}^{t+T}R(s_{i},a_{i})$ in task $e$, while negative skills $\pi_{c}^{-}$ encompass all the other sub-optimal skills with lower returns in task $e$.
-Thus, we can simply sample the trajectory with the ranked highest return from the replay buffer as the positive sample $\tau_{c}^{+}$, and the one with the lowest return as the negative sample $\tau_{c}^{-}$.
-The SaNCE loss is then minimised to bring the context embeddings of the highest return trajectories closer in the embedding space while distancing those of negative trajectories.
-Note that, at the end of the training, the top-ranked trajectories in the ranked replay buffer correspond to positive samples $\tau_{c}^{+}$ with high returns, and the lower-ranked ones are negative samples $\tau_{c}^{-}$ with low returns.
-However, before the agent is able to achieve high returns, all trajectories are with low returns.
-Therefore, our SaNCE loss is a soft version of the $K$-sample SaNCE estimator:
+SaNCE can be used alone or combined with other optimisation objectives to train context encoders in Meta-RL algorithms.
+For instance, integrating SaNCE with InfoNCE diversifies the negative sample space, with $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M$.
+The sample space for $I_{\text{Sa+InfoNCE}}$ is depicted in Figure 4(c) and further analysed in detail in Appendix C.
+4.4 Skill-aware trajectory sampling strategy
+In this section, we propose a practical trajectory sampling method.
+Methods focusing on skill diversity often rely heavily on accurately defining and identifying individual skills [Eysenbach et al., 2018].
+Some of these methods require a prior skill distribution, which is often inaccessible [Shi et al., 2022], and it is impractical to enumerate all possible skills that we hope the model to learn.
+Besides, we believe that distinctiveness of skills is inherently difficult to achieve — a slight difference in states can make two skills distinguishable, and not necessarily in a semantically meaningful way.
+Consequently, we do not directly teach any of these skills or assume any prior skill distribution.
+Diverse skills naturally emerge from the incentives of the SaMI learning objective in a multi-task setting, driven by the inherent need to develop generalisable skills.
+For example, in high-friction tasks, the agent must acquire the Pick&Place skill to avoid large frictional forces, whereas in high-mass tasks, the agent must learn the Push skill since it cannot lift the cube.
+In each task, we only identify whether the skills are optimal; in this way, under a multi-task setting, the agent will acquire a set of general skills that are applicable to many tasks.
+In a given task $e$, positive skills $\pi_{c}^{+}$ are defined as optimal skills achieving highest return $\sum_{i=t}^{t+T}R(s_{i},a_{i})$, whereas negative skills $\pi_{c}^{-}$ are those that result in lower returns.
+As a result, the positive sample $\tau_{c}^{+}$ consists of trajectories generated by the skills with the highest ranked returns, while the negative samples correspond to those with the lowest returns.
+This straightforward approach of selecting positive samples based on the ranked highest return effectively aligns with positive skills and mitigates the challenge of hard negative examples [Robinson et al., 2021].
+The SaNCE loss is then minimised to bring the context embeddings of the highest return trajectories closer while distancing those of negative trajectories.
+By the end of training, the top-ranked trajectories in the ranked replay buffer correspond to positive samples $\tau_{c}^{+}$ with high returns, while the lower-ranked trajectories represent negative samples $\tau_{c}^{-}$ with low returns.
+However, at the beginning of training, it is likely that all trajectories have low returns.
+Therefore, our SaNCE loss is a soft variant of the $K$-sample SaNCE:
 $\mathcal{L}_{\text{SaNCE}}=-\max\left(||\psi(\tau_{c}^{+}),\psi(\tau_{c}^{-})||_{L2},1\right)\cdot I_{\text{SaNCE}}$ (4)
@@
 5 Experiments
-We use a three-step process to demonstrate the benefits of SaMI in each environment and answer three questions: (1) Does optimising SaMI lead to increased returns during training and zero-shot generalisation (see Table 5.1 and 2)?; (2) Does SaMI help the RL agents to be versatile and embody multiple skills (see Figure 6)?; (3) Can SaNCE overcome the $\log$-$K$ curse in sample-limited scenarios (see Table 5.1 and 2, and Section 5.4)?
+Our experiments aim to answer the following questions: (1) Does optimising SaMI lead to increased returns during training and zero-shot generalisation (see Table 5.1 and 2)?; (2) Does SaMI help the RL agents to be versatile and embody multiple skills (see Figure 6)?; (3) Can SaNCE overcome the $\log$-$K$ curse in sample-limited scenarios (see Table 5.1 and 2, and Section 5.4)?
+Figure 6: (a) UMAP visualisation of context embeddings for the SaCCM in the Panda-gym environment, with points in the yellow box representing the Push skill in high-mass tasks.
+Heatmap of (b) success rate, (c) Push skill probability, and (d) Pick&Place skill probability for SaCCM.
+In large-mass scenarios, the Push skill is more likely to be executed than Pick&Place.
 5.1 Experimental setup
 Modified benchmarks with multiple environmental features.
-[^1] Our modified benchmarks are open-sourced at https://github.com/uoe-agents/Skill-aware-Panda-gym We demonstrate the efficacy of our method using two benchmarks, Panda-gym [Gallouédec et al., 2021] and MuJoCo [Todorov et al., 2012] (details in Sections 5.2 and 5.3).
-The benchmarks are modified to be influenced by multiple environmental features simultaneously.
-Environmental features are sampled at the start of each episode during both training and testing phases.
-During training, we uniform-randomly select a combination of environmental features from a training set.
-At test time, we evaluate each algorithm in unseen tasks with different environmental features outside the training range.
+[^1] Our modified benchmarks are open-sourced at https://github.com/uoe-agents/Skill-aware-Panda-gym We evaluate our method on two benchmarks, Panda-gym [Gallouédec et al., 2021] and MuJoCo [Todorov et al., 2012] (details in Sections 5.2 and 5.3).
+The benchmarks are modified to be influenced by multiple environmental features, which are sampled at the start of each episode during meta-training and meta-testing.
+During meta-training, we uniform-randomly select a combination of environmental features from a training task set.
+At test time, we evaluate each algorithm in unseen tasks with environmental features outside the training range.
 Generalisation performance is measured in two different regimes: moderate and extreme.
 The moderate regime draws environmental features from a closer range to the training range compared to the extreme.
-For all our experiments, we report the mean and standard deviation of the models trained over five seeds in both training and test tasks.
-Further experimental details are available in Appendix D.
+Our results report the mean and standard deviation of models trained over five seeds in both training and test tasks.
+Further details are available in Appendix D.
 Baselines.
-In our experiments, we primarily compare InfoNCE with SaNCE to demonstrate the performance improvements brought by SaNCE.
-Therefore, we compare our method with three prevailing and competitive baselines.
-First, we consider CCM [Fu et al., 2021], which is equipped with InfoNCE.
-Additionally, we consider TESAC [Yu et al., 2020], which employs a value function loss, allowing us to evaluate the impact on the context encoder without contrastive loss.
-Given that CCM and TESAC are using RNN encoder, we also consider PEARL [Rakelly et al., 2019], which utilises an MLP context encoder and a similar loss as TESAC.
-Our methods.
-[^2] Our code and experimental data are available at https://github.com/uoe-agents/SaMI We employ Soft Actor-Critic (SAC) [Haarnoja et al., 2018] as the base RL algorithm.
-SaNCE can be integrated with any Meta-RL algorithm.
-Hence, we equip two RL algorithms with SaNCE: (1) SaTESAC is TESAC with SaNCE, which solely uses SaNCE for contrastive learning, with a significantly smaller sample space than that of other algorithms, as shown in Figure 4(b); (2) SaCCM is CCM with SaNCE, where the contrastive learning combines InfoNCE and SaNCE, as shown in Figure 4(c).
-The purpose is to compare both the impact of different negative sample spaces on the contrastive learning and downstream RL tasks, and the adaptability of SaNCE across different algorithms.
-In Appendix G, we provide a comparison in MuJoCo with numerical results from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] because of the exact same environmental setting.
-The sample spaces of PEARL, TESAC, and CCM are all $|c|$ times larger than that of SaTESAC.
+The loss function of the context encoder in all algorithms consists of two key components: the RL loss and the contrastive loss.
+The RL loss $\mathcal{L}_{RL}$, which is the same across all methods, corresponds to the RL value function loss.
+Our primary comparison focuses on the contrastive loss, taking the form of either SaNCE, InfoNCE, or no contrastive loss.
+Accordingly, we select baselines based on contrastive loss: CCM [Fu et al., 2021], which utilises InfoNCE, and TESAC [Yu et al., 2020], which relies solely on the RL loss, allowing assessment of the context encoder without contrastive loss.
+Since CCM and TESAC use an RNN encoder, we also include PEARL [Rakelly et al., 2019], which utilises an MLP context encoder and follows the similar RL loss.
+Additionally, Appendix G includes comparisons with DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020], using the same environmental setup in the MuJoCo benchmark.
+Our method.
+[^2] Our code, video demos and experimental data are available at https://github.com/uoe-agents/SaMI We use Soft Actor-Critic (SAC) [Haarnoja et al., 2018] as the base RL algorithm, training agents for 1.6 million timesteps in each environment (details in Appendix D.3).
+SaNCE is a simple objective based on MI that can be used to train any context encoder.
+We integrate SaNCE into two Meta-RL algorithms: (1) SaTESAC is TESAC with SaNCE, which uses SaNCE for contrastive learning, using a $|c|$ times smaller sample space (Figure 4(b)); (2) SaCCM is CCM with SaNCE, where the contrastive learning combines InfoNCE and SaNCE, as shown in Figure 4(c).
 5.2 Panda-gym
 Task description.
-Our modified Panda-gym benchmark contains a robot arm control task using the Franka Emika Panda [Gallouédec et al., 2021], where the robot needs to move a cube to a target position.
-Unlike previous works, the robot can flexibly execute different skills (Push and Pick&Place) to solve tasks.
-In our experiments, we simultaneously modify multiple environmental features (cube mass and table friction) that characterise the transition dynamics.
-In some environmental settings,
-Table 1: Comparison of success rate with baselines in Panda-gym (over 5 seeds).The blue background indicates that our method surpasses PEARL, TESAC and CCM.
+Our modified Panda-gym benchmark involves a robot arm control task using the Franka Emika Panda [Gallouédec et al., 2021], where the robot moves a cube to a target position.
+Unlike previous works, we simultaneously modify multiple environmental features (cube mass and table friction) that characterise the transition dynamics, and the robot can flexibly execute different skills (Push and Pick&Place) for different tasks.
+This environment requires high skill diversity; for instance, the agent must use Pick&Place in high-friction tasks and Push in high-mass tasks.
+Table 1: Comparison of success rate $\pm$ standard deviation with baselines in Panda-gym (over 5 seeds).
 Bold text signifies the highest average return.
+* next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI.
+All significance claims based on paried t-tests with significance threshold of $p<0.05$.
 Training Test (moderate) Test (extreme)
 PEARL 0.42$\pm$0.19 0.10$\pm$0.06 0.11$\pm$0.05
-TESAC 0.64$\pm$0.16 0.33$\pm$0.23 0.27$\pm$0.17
-CCM 0.82$\pm$0.04 0.50$\pm$0.23 0.29$\pm$0.23
-SaTESAC 0.92$\pm$0.05 0.55$\pm$0.25 0.40$\pm$0.35
-SaCCM 0.94$\pm$0.02 0.58$\pm$0.27 0.46$\pm$0.36
-only a specific skill can accomplish the task. For example, an agent needs to learn the Pick&Place skill to avoid high friction. Thus, the agent must determine when and which skill to execute.
-Evaluating zero-shot generalisation.
+TESAC 0.50$\pm$0.22 0.31$\pm$0.20 0.22$\pm$0.21
+CCM 0.80$\pm$0.19 0.49$\pm$0.23 0.29$\pm$0.28
+SaTESAC 0.92$\pm$0.04 0.56$\pm$0.24 0.37$\pm$0.34
+SaCCM 0.93$\pm$0.05 0.57$\pm$0.26 0.36$\pm$0.35
+Results and skill analysis.
 As shown in Table 5.1, SaTESAC and SaCCM achieve superior generalisation performance compared to PEARL, TESAC, and CCM, with a smaller sample space.
-In training, moderate test, and extreme test tasks, SaTESAC achieved up to 43.75%, 66.67%, and 48.15% higher success rates, respectively, compared to TESAC, while SaCCM improved success rates by 14.63%, 16.00%, and 58.62% compared to CCM.
-Figure 6: (a) t-SNE visualisation of context embeddings in the Panda-gym environment.
-The clustered points in the yellow box correspond to the Push skill executed in scenarios with a large mass.
-Heatmap of (b) success rate, and probability of learned (c) Push skill and (d) Pick&Place skill of SaCCM.
-Under large mass scenarios, the probability of executing Push skill is higher than Pick&Place skill.
-Skill analysis.
-We visualise the context embedding of SaCCM using t-SNE [Van der Maaten and Hinton, 2008].
-We determine the skills by detecting contact points between the end effector and the cube and between the cube and the table (see Appendix D for more details), and then employ Heatmaps [Waskom, 2021] to visualise skills executed by the agent.
-As shown in Figure 6, we found that when the cube mass is large (30 Kg and 10 Kg), the agent learned the Push skill (corresponding to the clustered points in the yellow bounding box in Figure 6(a).
-With smaller masses, the agent learned the Pick&Place skill.
-However, as shown in Figure 15 in Appendix F.1, CCM did not exhibit clear skill grouping.
-This indicates that SaMI extracts high-quality skill-related information from the trajectories and aids in embodying diverse skills.
-We provide additional visualisation based on both t-SNE and PCA [Jolliffe, 2002], and heatmap results in Appendix F.
+The t-test results in Table 5.1 show that SaMI significantly improves success rates across training, moderate, and extreme test sets at a 0.05 significance level.
+Video demos2 show that agents equipped with SaMI acquired multiple skills (Push, Pick&Place) to handle various tasks.
+When faced with an unknown task, the agents explore by attempting to lift the cube, infer the context, and adjust their skills accordingly within the episode.
+We visualised the context embeddings using UMAP [McInnes et al., 2020] (Figure 6(a) and Appendix F) and t-SNE [Van der Maaten and Hinton, 2008] (Appendix F), plotting the final step from 100 tests per task.
+Skills were identified through contact points between the end effector, cube, and table (see Appendix D for more details), and heatmaps [Waskom, 2021] were used to visualise executed skills.
+Figure 6 shows that SaCCM agents learned the Push skill for large cube masses (30 Kg, 10 Kg) and Pick&Place for smaller masses, while CCM showed no clear skill grouping (Figure 16 in Appendix F.1).
+Overall, SaMI incentivises agents to autonomously learn diverse skills, enhancing generalisation across a wider range of tasks.
+Specifically, through the cycle of effective exploration, context inference, and adaptation, diverse skills emerge solely from the data.
+Further visualisation results are in Appendix F.
+Table 2: Comparison of average return $\pm$ standard deviation with baselines in modified MuJoCo benchmark (over 5 seeds).
+Bold number signifies the highest return.
+* next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI.
+All significance claims based on t-tests with significance threshold of $p<0.05$.
+Crippled Ant Crippled Half-cheetah
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+PEARL 1682$\pm$73 996$\pm$21 888$\pm$31 1998$\pm$973 698$\pm$548 746$\pm$1092
+TESAC 2139$\pm$90 1952$\pm$40 1048$\pm$124 3967$\pm$955 874$\pm$901 846$\pm$849
+CCM 2361$\pm$114 2047$\pm$83 1527$\pm$301 3481$\pm$488 821$\pm$575 873$\pm$914
+SaTESAC 2638$\pm$406 2379$\pm$528 2131$\pm$132 4328$\pm$1092 1143$\pm$664 1540$\pm$1094
+SaCCM 2355$\pm$170 2310$\pm$314 2007$\pm$68 4478$\pm$1131 1007$\pm$568 1027$\pm$782
+Ant Half-cheetah
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+PEARL 5153$\pm$581 3873$\pm$235 3802$\pm$409 5802$\pm$773 2190$\pm$970 1346$\pm$692
+TESAC 6789$\pm$451 4705$\pm$279 4108$\pm$369 6298$\pm$2310 3173$\pm$1210 1159$\pm$338
+CCM 6901$\pm$567 5179$\pm$902 4700$\pm$696 6955$\pm$788 3963$\pm$622 1325$\pm$269
+SaTESAC 7314$\pm$545 5513$\pm$648 4940$\pm$531 7430$\pm$1026 4058$\pm$890 1780$\pm$102
+SaCCM 7478$\pm$539 5717$\pm$488 5215$\pm$377 7154$\pm$965 3849$\pm$689 1926$\pm$218
+SlimHumanoid HumanoidStandup
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+PEARL 6947$\pm$3541 3697$\pm$2674 2018$\pm$907 95456$\pm$13445 63242$\pm$13546 64224$\pm$15467
+TESAC 8437$\pm$1798 6989$\pm$1301 3760$\pm$308 158384$\pm$14455 153944$\pm$15046 74220$\pm$19980
+CCM 7696$\pm$1907 5784$\pm$531 2887$\pm$1058 146480$\pm$33745 154601$\pm$16291 94991$\pm$15258
+SaTESAC 10216$\pm$1620 7886$\pm$2203 6123$\pm$1403 178142$\pm$10081 168337$\pm$12123 133335$\pm$24607
+SaCCM 9312$\pm$705 7430$\pm$1587 6473$\pm$2001 187930$\pm$19338 181033$\pm$14628 141750$\pm$27426
+Hopper Crippled Hopper
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+PEARL 934$\pm$242 874$\pm$366 799$\pm$298 3091$\pm$298 2387$\pm$656 456$\pm$235
+TESAC 1492$\pm$59 1499$\pm$35 1459$\pm$72 3575$\pm$192 3298$\pm$551 722$\pm$161
+CCM 1484$\pm$54 1446$\pm$64 1452$\pm$58 3455$\pm$301 3409$\pm$239 1009$\pm$289
+SaTESAC 1502$\pm$20 1453$\pm$39 1447$\pm$14 3391$\pm$84 3262$\pm$166 1839$\pm$130
+SaCCM 1462$\pm$45 1462$\pm$14 1451$\pm$67 3449$\pm$103 3390$\pm$211 2059$\pm$221
+Walker Crippled Walker
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+PEARL 7524$\pm$2455 3355$\pm$2555 1984$\pm$356 7899$\pm$2532 4377$\pm$2563 2965$\pm$1426
+TESAC 7747$\pm$1772 4355$\pm$1530 2581$\pm$407 9908$\pm$1561 5929$\pm$1971 3041$\pm$912
+CCM 8136$\pm$557 5476$\pm$803 2519$\pm$682 10317$\pm$1137 6233$\pm$1869 3098$\pm$821
+SaTESAC 8675$\pm$752 5840$\pm$676 3632$\pm$404 10389$\pm$1031 8387$\pm$1291 4280$\pm$485
+SaCCM 8361$\pm$586 5779$\pm$691 3481$\pm$332 10496$\pm$951 8235$\pm$1212 4824$\pm$839
 5.3 MuJoCo
 Task description.
-The modified MuJoCo benchmark [Mu et al., 2022, Lee et al., 2020] contains 6 typical robotic control environments based on the MuJoCo physics engine [Todorov et al., 2012].
-Hopper, Half-cheetah, Ant, and SlimHumanoid are affected by multiple continuous environmental features (i.e., mass, damping) that characterise the transition dynamics.
-Both Crippled Ant and Crippled Half-cheetah are more difficult because they are also affected by a discrete environmental feature as one of the actuators in the legs is randomly crippled to change the transition dynamics.
-Table 2: Comparison of average return with baselines in modified MuJoCo benchmark (over 5 seeds).
-The blue background indicates that our method surpasses PEARL, TESAC and CCM.
-Bold text signifies the highest average return.
-Crippled Ant Crippled Half-cheetah
-Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
-PEARL 182$\pm$73 96$\pm$21 88$\pm$31 2538$\pm$783 1028$\pm$445 908$\pm$185
-TESAC 485$\pm$83 129$\pm$32 117$\pm$32 4512$\pm$737 1708$\pm$360 1550$\pm$274
-CCM 814$\pm$23 146$\pm$13 158$\pm$31 4153$\pm$569 1417$\pm$710 995$\pm$188
-SaTESAC 568$\pm$70 176$\pm$9 196$\pm$17 5380$\pm$917 2334$\pm$169 1848$\pm$341
-SaCCM 719$\pm$82 215$\pm$6 219$\pm$9 6210$\pm$402 2656$\pm$231 2070$\pm$651
-Ant Half-cheetah
-Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
-PEARL 153$\pm$63 73$\pm$25 39$\pm$40 1802$\pm$773 530$\pm$270 346$\pm$102
-TESAC 761$\pm$234 360$\pm$23 276$\pm$63 5817$\pm$3064 2843$\pm$1303 1074$\pm$446
-CCM 1268$\pm$362 1078$\pm$492 548$\pm$177 6839$\pm$542 3053$\pm$347 1422$\pm$332
-SaTESAC 911$\pm$54 590$\pm$111 552$\pm$119 7303$\pm$1353 4077$\pm$424 1723$\pm$88
-SaCCM 1015$\pm$60 629$\pm$126 588$\pm$92 7662$\pm$316 4384$\pm$223 1828$\pm$52
-SlimHumanoid Hopper
-Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
-PEARL 6947$\pm$3541 3697$\pm$2674 2018$\pm$907 934$\pm$242 874$\pm$366 799$\pm$298
-TESAC 9095$\pm$894 7180$\pm$711 3709$\pm$35 1530$\pm$39 1476$\pm$9 1491$\pm$79
-CCM 7297$\pm$790 5856$\pm$290 3501$\pm$877 1488$\pm$18 1420$\pm$46 1451$\pm$76
-SaTESAC 10501$\pm$2102 8483$\pm$1502 6669$\pm$883 1500$\pm$24 1452$\pm$36 1449$\pm$15
-SaCCM 9738$\pm$92 8389$\pm$231 5606$\pm$2027 1472$\pm$53 1456$\pm$13 1446$\pm$94
-Results.
-Table 2 shows the average return of our method and baselines on both training and test tasks.
-Incorporating SaNCE into TESAC resulted in SaTESAC’s average return increasing by 19.41% in training tasks across all environments, 39.70% in moderate test tasks, and 65.40% in extreme test tasks.
-Similarly, adding SaNCE to CCM led to SaCCM achieving up to 31.69%, 44.83%, and 40.88% higher average returns in training tasks, moderate test tasks, and extreme test tasks, respectively.
-Overall, SaTESAC and SaCCM outperform baselines in both training and testing in most of the tasks, in which the Ant, Crippled Ant and Hopper tasks are exceptions.
-Comparing the visualisation of context embeddings for CCM and SACCM in the Ant, Crippled Ant, and Hopper environments (Appendix F.2), introducing SaNCE does not modify the grouping of context embeddings across different tasks.
-Notably in Hopper environment, all algorithms maintain consistent average returns when zero-shot generalising to unseen test tasks, with no clear clustering of context embeddings across different tasks (Table 22 in Appendix F.2).
-This indicates that a single skill may generalise across numerous tasks within the Hopper environment.
-Hence, we can draw an insight that SaMI is more suitable for environments that require multiple skills to adapt to different environmental features.
-5.4 Analysis of the $\log$-$K$ curse in sample-limited scenarios
-In this section, we further analyse whether SaNCE can overcome the $\log$-$K$ curse.
-During the training phases, we sample the environmental features at the beginning of each episode.
-Therefore, throughout the training process, the context encoder needs to learn the context embedding distribution of multiple tasks.
-Because InfoNCE requires sampling negative samples across all tasks, and SaNCE samples negative samples from the current task, SaNCE’s negative sample space is $|c|$ times smaller than that of InfoNCE.
-For example, in the SlimHumanoid environment, where both mass and damping have five values, there can be a maximum of 25 different tasks, making the sampling space of InfoNCE potentially 25 times larger than that of SaNCE.
-According to Table 5.1 and 2, RL algorithms equipped with SaNCE can achieve better or comparable performance with a much smaller number of negative samples $K$ than InfoNCE.
-This indicates that SaNCE indeed plays a significant role in addressing the $\log$-$K$ curse, and the SaMI objective helps the contrastive context encoder extract information that is crucial for downstream RL tasks.
-Furthermore, the number of negative samples $K$ also relates to two hyperparameters: buffer size, which directly determines the size of the negative sample space; and contrastive batch size, which determines the number of samples used for training contrastive context encoder at each update step.
-Therefore, we conducted further analysis on these two hyperparameters.
-As shown in Figure 7, we find that reductions in buffer size and contrastive batch size do not significantly decrease the average return for SaCCM and SaTESAC, which exhibit state-of-the-art performance with small buffers and contrastive batch sizes.
-Note that the results in Table 2 correspond to a buffer size of 100,000 and a contrastive batch size of 12.
-Experimental results for all environments (Appendix E.2) further demonstrate that SaNCE is not highly sensitive to changes in $K$, and indeed shows great potential in overcoming the $\log$-$K$ curse.
+We extended the modified MuJoCo benchmark introduced in DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020].
+It contains ten typical robotic control environments based on the MuJoCo physics engine [Todorov et al., 2012].
+Hopper, Walker, Half-cheetah, Ant, HumanoidStandup, and SlimHumanoid are influenced by continuous environmental features (i.e., mass, damping) that affect transition dynamics.
+Crippled Ant, Crippled Hopper, Crippled Walker, and Crippled Half-cheetah are more challenging due to the addition of discrete environmental features (i.e., randomly crippled leg joints), requiring agents to master different skills (e.g., switching from running to crawling after a leg is crippled).
+Results and skill analysis.
+Table 2 shows the average return of our method and baselines on training and test tasks.
+SaTESAC and SaCCM achieved higher returns in most tasks, except for Ant, Half-Cheetah, and Hopper, where only a single skill was needed.
+For instance, the Hopper robot learned to hop forward, adapting to different mass values.
+When environments become complex and require diverse skills for different tasks (Crippled Ant, Crippled Hopper, Crippled Half-Cheetah, SlimHumanoid, HumanoidStandup, and Crippled Walker), SaNCE brings significant improvements.
+For example, when the Ant robot has 3 or 4 legs available, it learns to roll to generalise across varying mass and damping.
+In more challenging zero-shot settings, when only 2 legs are available, the ant robot can no longer roll and it adapts by walking using its 2 healthy legs.
+This aligns with the results in Table 2, where SaMI significantly improved performance in extreme test sets.
+In summary, i) SaMI helps the RL agents to be versatile and embody multiple skills; ii) SaMI leads to increased returns during training and zero-shot generalisation, especially in environments that require different skills.
+Our video demos2 and visualisation results in Appendix F.2 show different skills in all environments.
+5.4 Analysis of the $\boldsymbol{\log}$-$\boldsymbol{K}$ curse in sample-limited scenarios
+This section analyses whether SaNCE can overcome the $\log$-$K$ curse.
+During training, environmental features are sampled at the start of each episode, requiring the context encoder to learn the context embedding distribution across multiple tasks.
+Since InfoNCE samples negative examples from all tasks and SaNCE samples from the current task, SaNCE’s negative sample space is $|c|$ times smaller than InfoNCE’s.
+For instance, in the SlimHumanoid environment, where both mass and damping have five values, InfoNCE’s sampling space can be 25 times larger than SaNCE’s.
+As shown in Tables 5.1 and 2, RL algorithms using SaNCE achieve better or comparable performance with significantly fewer negative samples ($K$) than InfoNCE.
+This suggests SaNCE effectively addresses the $\log$-$K$ curse, and the SaMI objective helps the contrastive context encoder extract critical information for downstream RL tasks.
+The number of negative samples $K$ is influenced by two hyperparameters: buffer size, which determines the negative sample space, and contrastive batch size, which controls the number of samples used to train the contrastive context encoder per update.
+We analysed these hyperparameters further, and as shown in Figure 7, reductions in buffer and contrastive batch size do not significantly impact the average return for SaCCM and SaTESAC, which maintain state-of-the-art performance with small buffers and batch sizes.
+The results in Table 2 correspond to a buffer size of 100,000 and a batch size of 12.
+Results across all environments (refer to Appendix E.2) show that SaNCE exhibits low sensitivity to $K$, highlighting its potential to overcome the $\log$-$K$ curse.
+Figure 7: Effect of (a) buffer size (TESAC, CCM, SaTESAC, SaCCM) and (b) contrastive batch size (CCM, SaTESAC, SaCCM) in the SlimHumanoid environment.
 6 Conclusion and future work
-In this paper, we propose a Skill-aware Mutual Information (SaMI) to learn context embeddings for zero-shot generalisation in downstream RL tasks, and a Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI and overcome the $\log$-$K$ curse.
-RL algorithms equipped with SaMI have achieved state-of-the-art performance in the MuJoCo and Panda-gym benchmarks.
-Through visual analysis of context embeddings and heatmap analysis of learned skills in Panda-gym, we confirm that SaMI assists the RL agent in embodying diverse skills.
-These experimental observations indicate that the context encoder, learned by maximising SaMI, helps the RL agent generalise across various tasks by executing different skills.
-Notably, the optimisation process of SaNCE utilises a far smaller negative sample space than baselines.
-Coupled with further experimental analysis of buffer size and contrastive batch size on MuJoCo, we demonstrate that SaNCE helps overcome the $\log$-$K$ curse.
-Figure 7: Changes of average return in the SlimHumanoid environment.
-(a) The impact of buffer size on RL control performance (i.e., TESAC, CCM, SaTESAC, SaCCM).
-(b) The impact of contrastive batch size on the contrastive context encoder (i.e., CCM, SaTESAC, SaCCM).
-Our results indicate the importance of aligning the optimisation objectives of representation learning and downstream optimal control to enhance task performance and improve data efficiency.
+Zero-shot generalisation has been a longstanding challenge concerning Meta-RL agents, with skill diversity and sample efficiency being key to generalising to previously unseen environments.
+In this paper, we proposed Skill-aware Mutual Information (SaMI) to learn context embeddings for zero-shot generalisation in downstream RL tasks, and Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI and overcome the $\log$-$K$ curse, along with a practical skill-aware trajectory sampling strategy.
+Experimental results showed that RL algorithms equipped with SaMI achieved state-of-the-art performance in MuJoCo and Panda-gym benchmarks, particularly in zero-shot generalisation within more complex environments.
+During the zero-shot generalisation, when faced with an unseen task, SaMI assists agents in exploring effectively, inferring context, and rapidly adapting their skills within the current episode.
+SaNCE’s optimisation uses a significantly smaller negative sample space than baselines, and our analysis on buffer and contrastive batch sizes demonstrated its effectiveness in addressing the $\log$-$K$ curse.
 Given that environmental features are often interdependent, such as a cube’s material correlating with friction and mass, SaMI does not introduce independence assumptions like DOMINO [Mu et al., 2022].
-Therefore, future work will focus on verifying and enhancing SaMI’s potential in more complex tasks where environmental features are correlated.
+Therefore, future work will focus on verifying and enhancing SaMI’s potential in more complex tasks where environmental features are correlated [Dunion et al., 2023a].
 This will contribute to our ultimate goal: developing a generalist and versatile agent capable of working across multiple tasks and even real-world tasks in the near future.
 Appendix A Proof of Lemma 1
-Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, A $K$-sample InfoNCE estimator is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$.
-InfoNCE is obtained by comparing positive pairs $(x,y_{1})$ and negative pairs $(x,y_{k})$, where $y_{k}\sim y_{2:K}$, as follows:
+Given a query $x$ and a set $Y=\{y_{1},\dots,y_{K}\}$ of $K$ random samples, containing one positive sample $y_{1}$ and $K-1$ negative samples drawn from the distribution $p(y)$, a $K$-sample InfoNCE estimator is obtained by comparing pairs sampled from the joint distribution $(x,y_{1})\sim p(x,y)$ with pairs $(x,y_{k})$, constructed using the set of negative examples $y_{2:K}$.
+InfoNCE compares the positive pairs $(x,y_{1})$ with the negative pairs $(x,y_{k})$, where $y_{k}\sim y_{2:K}$, as follows:
 $I_{\text{InfoNCE}}(x;y|\psi,K)$ $=\mathbb{E}_{p(x,y_{1})p(y_{2:K})}\left[\log\left(\frac{f_{\psi}(x,y_{1})}{\frac{1}{K}\sum_{k=1}^{K}f_{\psi}(x,y_{k})}\right)\right]$ (5)
 Step 1.
-Let us prove the $K$-sample InfoNCE estimator is upper bounded by $\log K$.
+Let us prove that the $K$-sample InfoNCE estimator is upper-bounded by $\log K$.
 According to Mu et al.
@@
 Hence, we have $I_{\text{InfoNCE}}(x;y|\psi,K)\leq\log K$.
-Step 2. We have the $I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K)$ which are rightmost and the leftmost side in Lemma 1 according to:
+Step 2. We have the $I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K)$ according to:
 Proposition 1
-[Poole et al., 2019] A $K$-sample estimator is an asymptotically tight lower bound to the MI, i.e.,
+[Poole et al., 2019] A $K$-sample estimator is an asymptotically tight lower bound on the MI, i.e.,
 $I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K),\lim_{x\to+\infty}I_{\text{InfoNCE}}(x;y|\psi,K)\to I(x;y)$
@@
 In this research, the context encoder $\psi$ in $f_{\psi}(x,y)$ is implemented using an RNN to approximate $\frac{p(y|x)}{p(y)}$ [Oord et al., 2019].
-For most deep learning platforms, with a powerful learner for $\psi$ and a limited size $K$ such that $I(x;y)\geq\log K$, we can reasonably expect $I_{\text{InfoNCE}}\approx\log K$ after a few training epochs.
+With a sufficiently powerful deep learning model for $\psi$ and a finite sample size $K$, such that $I(x;y)\geq\log K$, we can reasonably expect that $I_{\text{InfoNCE}}\approx\log K$ after a few training epochs.
 Therefore, during training, when $K\ll+\infty$, we always have $I(x;y)\geq\log K$.
@@
 Let us prove that the $K$-sample InfoNCE bound is asymptotically tight.
-The specific choice of context encoder $\psi$ relates to the tightness of the $K$-sample NCE bound.
-InfoNCE [Oord et al., 2019] sets $f_{\psi}(x,y)\propto\frac{p(y|x)}{p(y)}$ to model a density ratio which preserves the MI between $x$ and $y$, where $\propto$ stands for ‘proportional to’ (i.e.
-up to a multiplicative constant).
-Let us plug in $f_{\psi}(x,y)=f_{\psi}^{*}(x,y)=\frac{p(y|x)}{p(y)}$ into InfoNCE, and we have
+The specific choice of the context encoder $\psi$ influences the tightness of the $K$-sample NCE bound.
+InfoNCE [Oord et al., 2019] sets $f_{\psi}(x,y)\propto\frac{p(y|x)}{p(y)}$ to model a density ratio that preserves the MI between $x$ and $y$, where $\propto$ stands for ’proportional to’ (i.e., up to a multiplicative constant).
+Substituting $f_{\psi}(x,y)=f_{\psi}^{*}(x,y)=\frac{p(y|x)}{p(y)}$ into InfoNCE, we obtain:
 $I_{\text{InfoNCE}}(x;y|\psi,K)$ $=\mathbb{E}\left[\log\left(\frac{f_{\psi}^{*}(x,y_{1})}{\sum_{k=1}^{K}f_{\psi}^{*}(x,y_{k})}\right)\right]+\log K$
@@
 Now taking $K\to+\infty$, the last two terms cancel out.
-Putting it together. Combining $I(x;y)\geq\log K$ with Proposition 1 and Equation 6, we have Lemma 1:
+Putting it together. Combining $I(x;y)\geq\log K$ with Proposition 1 and Equation (6), we have Lemma 1:
 ${I}_{\text{InfoNCE}}(x;y|\psi,K)\leq\log K\leq I(x;y).$ (8)
-Besides, according to Equation 7, with sample size $K\to+\infty$, $K$-sample InfoNCE bound is sharp and approaches to the true MI $I(x;y)$, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K\approx I(x;y)$.
+Moreover, according to Equation (7), as the sample size $K\to+\infty$, the $K$-sample InfoNCE bound becomes sharp and approaches the true MI $I(x;y)$, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K\approx I(x;y)$.
 Appendix B Proof for Lemma 2
 Step 1.
-According to Lemma 1, we have ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$ (shown in Figure 3).
+According to Lemma 1, we have ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$ (shown in Figure 3).
 Step 2.
-Let us prove SaNCE is a $K$-sample SaNCE estimator and upper bounded by ${\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$.
-Because $\frac{f_{\psi}(c,\pi_{c},\tau_{c}^{+})}{f_{\psi}(c,\pi_{c},\tau_{c}^{+})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c,k}^{-})}\leq 1$ [Mu et al., 2022], so we have:
-$\quad{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
+Let us prove that SaNCE is a $K$-sample SaNCE estimator and is upper bounded by ${\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$.
+Since $\frac{f_{\psi}(c,\pi_{c},\tau_{c}^{+})}{f_{\psi}(c,\pi_{c},\tau_{c}^{+})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c,k}^{-})}\leq 1$ [Mu et al., 2022], we have:
+$\quad{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
 $=\mathbb{E}_{p(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K}^{-})}\left[\log\left(\frac{K\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
@@
 $\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$ (9)
-Hence, we have ${\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$ (similar to Equation 6).
+Thus, we obtain ${\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$, similar to Equation (6).
 Step 3.
-With the definition of $K^{*}$, we can prove ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
-In task $e_{1}$ with context embedding $c_{1}$, SaNCE obtains positive and negative samples from the current task $e_{1}$.
-Hence the variable $c=c_{1}$ is constant, we have:
-$\quad{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
+With the definition of $K^{*}$, we can prove that ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
+In task $e_{1}$, SaNCE obtains positive and negative samples from the current task $e_{1}$.
+Since the variable $c=c_{1}$ is constant, we have:
+$\quad{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
 $=\mathbb{E}_{p(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})p(\tau_{c_{1},2:K}^{-})}\left[\log\left(\frac{K\cdot f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})}{f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1}}^{+})+\sum_{k=2}^{K}f_{\psi}(c_{1},\pi_{c_{1}},\tau_{c_{1},k}^{-})}\right)\right]$
@@
 $\approx{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K^{*}_{\text{SaNCE}}}$ (10)
-The required sample size $K^{*}_{\text{SaNCE}}=|c_{1}|\cdot|\pi|\cdot M=|\pi|\cdot M$.
-With $K\to K^{*}_{\text{SaNCE}}$, ${\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$.
-Correspondingly, for InfoNCE, in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{1}$ generated after executing the skill $\pi_{1}$ in task $e_{1}$, and negative samples are trajectories $\{\tau_{c_{2}}^{-},...\}$ from other tasks $\{e_{2},...\}$.
+The required sample size is $K^{*}_{\text{SaNCE}}=|c_{1}|\cdot|\pi|\cdot M=|\pi|\cdot M$.
+As $K\to K^{*}_{\text{SaNCE}}$, we have ${\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$.
+Correspondingly, for InfoNCE, in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{1}$ generated after executing the skill $\pi_{1}$ in task $e_{1}$, while negative samples are trajectories $\{\tau_{c_{2}}^{-},...\}$ from other tasks $\{e_{2},...\}$.
 Under the definition of $K^{*}$, we have:
-$\quad{\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
+$\quad{\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
 $=\mathbb{E}_{p(c,\pi_{c},\tau_{c_{1}})p({\tau_{c_{2:|c|},2:K}})}\left[\log\left(\frac{K\cdot f_{\psi}(c,\pi_{c},\tau_{c_{1}})}{f_{\psi}(c,\pi_{c},\tau_{c_{1}})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c_{2:|c|},k})}\right)\right]$
@@
 where $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M\approx|c|\cdot K^{*}_{\text{SaNCE}}$.
-In real-world robotic control tasks, the sample space size significantly increases due to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$.
-$|c|$ is the number of different tasks and increases exponentially due to the permutations and combinations of $N$ environmental features.
-When current task $e_{1}$ has context embedding $c_{1}$, the $c_{2:|c|}$ refer to the context embedding for the other tasks.
-With $K\to K^{*}_{\text{InfoNCE}}$, ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$.
-Hence, during the training process, ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
+In real-world robotic control tasks, the sample space size increases significantly due to multiple environmental features $e=\{e^{0},e^{1},\dots,e^{N}\}$.
+The number of different tasks $|c|$ grows exponentially due to the permutations and combinations of the $N$ environmental features.
+When the current task $e_{1}$ has context embedding $c_{1}$, the $c_{2:|c|}$ refer to the context embeddings for the other tasks.
+As $K\to K^{*}_{\text{InfoNCE}}$, we have ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$.
+Thus, during the training process, ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
 Step 4.
-What remains is to show ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$.
-The MI for three variables is also called interaction information.
-According to the definition in McGill [1954], the SaMI can be presented as:
-Proposition 2
-For the case of three variables, the MI could be written as ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}={\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}-I(c;\tau_{c}|\pi_{c})$.
-Using this proposition, one can see that in the case of three variables, interaction information quantifies how much the information shared between two variables differs from what they share if the third variable is known.
-Several properties of interaction information in the case of three variables have been studied in the literature.
-Specifically, Yeung [1991] showed that
-$-\min\{I(x;\tau_{c}|\pi_{c}),I(c;\pi_{c}|\tau_{c}),I(\tau_{c};\pi_{c}|c)\}\leq I(c;\tau_{c};\pi_{c})\leq\min\{I(c;\tau_{c}),I(c;\pi_{c}),I(\tau_{c};\pi_{c})\}$ (12)
-Combining Equation 12 and Proposition 2, we have ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$.
+According to the definition of SaMI in Equation (2), we have ${\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.265625,0.5234375,0.78125}I(c;\tau_{c})}$, as illustrated using Venn diagrams in Figure 8 (a) and (b).
+SaMI is formulated based on interaction information [McGill, 1954], which aims to capture the relationships among multivariate variables by quantifying the amount of information (redundancy or synergy) shared among three variables.
+Interaction information has been less extensively studied, partly due to its challenging interpretation from both information theory and neuroscience perspectives, as it can be either positive or negative [Li, 2022].
+Figure 8: Venn diagrams illustrating (a) mutual information $I(c;\tau_{c})$, (b) interaction information $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$, and (c) the MDP graph of the context embedding $c$, skill $\pi_{c}$, and trajectory $\tau_{c}$, which represents a common-cause structure [Neuberg, 2003].
+In the Meta-RL setting, the MDP (causal) graph structure, shown in Figure 8 (c), illustrates that the context embedding $c$, skill $\pi_{c}$, and trajectory $\tau_{c}$ form a common-cause structure, where $c$ acts as a shared cause influencing both the skill $\pi_{c}$ and the trajectory $\tau_{c}$.
+Consequently, $I(\pi_{c},\tau_{c}\mid c)<I(\pi_{c},\tau_{c})$.
+Therefore, $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})>0$ is guaranteed, making it interpretable in real-world robotic control tasks.
 Putting it together.
-Hence, we have Lemma 2: we always have ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$ (shown in Figure 3), while learning a skill-aware context encoder $\psi$ with SaNCE estimator.
-$K^{*}_{\text{SaNCE}}<<K^{*}_{\text{InfoNCE}}$ so that $I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$ is a much tighter lower bound of the true $I_{\text{SaMI}(c;\pi_{c};\tau_{c})}$ than ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$.
+Thus, we establish Lemma 2: we always have ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.265625,0.5234375,0.78125}I(c;\tau_{c})}$, as shown in Figure 3, while learning a skill-aware context encoder $\psi$ with the SaNCE estimator.
+Since $K^{*}_{\text{SaNCE}}\ll K^{*}_{\text{InfoNCE}}$, ${\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ serves as a much tighter lower bound for the true $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ than ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$.
 Appendix C Sample size of $I_{\text{Sa+InfoNCE}}$
-We illustrate the sample size of $I_{\text{Sa+InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$ in this section.
-Sa+InfoNCE plugs SaNCE into InfoNCE, and has positive samples $\tau_{c_{1}}^{+}$ from task $e_{1}$ after executing skill $\pi_{c_{1}}^{+}$, and negative samples are trajectories $\tau_{c_{1:K}}^{-}$ from executing skills $\pi_{c_{1:K}}^{-}$ in task $e_{1:K}$, respectively.
-Therefore, this is equivalent to observing the variable $c$, then observing the variable $\pi_{c}$, i.e., sampling from the distribution $p(\pi_{c},\tau_{c}|c)p(c)$.
+In this section, we illustrate the sample size of $I_{\text{Sa+InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$.
+Sa+InfoNCE incorporates SaNCE into InfoNCE, using positive samples $\tau_{c_{1}}^{+}$ from task $e_{1}$ after executing skill $\pi_{c_{1}}^{+}$, and negative samples are trajectories $\tau_{c_{1:K}}^{-}$ from executing skills $\pi_{c_{1:K}}^{-}$ in tasks $e_{1:K}$, respectively.
+Therefore, this approach is equivalent to first observing the variable $c$ and then observing the variable $\pi_{c}$, i.e., sampling from the distribution $p(\pi_{c},\tau_{c}|c)p(c)$.
 We have:
@@
 $\approx{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K^{*}_{\text{Sa+InfoNCE}}}$ (12)
-It should be noted that such a combination will increase the size of the negative sample space, i.e., $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M\geq K^{*}_{\text{SaNCE}}$.
-Therefore, with the same number $K$ of samples, $I_{\text{Sa+InfoNCE}}$ is less precise and looser than $I_{\text{SaNCE}}$.
-Hence, a trade-off between sample diversity and the precision of the $K$-sample estimator is required.
+It should be noted that such a combination increases the size of the negative sample space, i.e., $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M\geq K^{*}_{\text{SaNCE}}$.
+The misaligned bars in Figure 4 (c) illustrate that negative sample spaces may vary across tasks.
+This variation arises because we define negative samples as trajectories with low returns, making the size of the negative sample space influenced by sampling randomness.
+With the same number $K$ of samples, $I_{\text{Sa+InfoNCE}}$ is less precise and looser than $I_{\text{SaNCE}}$.
+Therefore, a trade-off between sample diversity and the precision of the $K$-sample estimator is required.
 Appendix D Environmental setup
-Figure 8: (a) Modified Panda-gym benchmarks, (b) the training tasks, the moderate test tasks and the extreme test tasks.
-The moderate test task setting just has a combinatorial interpolation.
-The extreme test task setting has unseen ranges of environmental features and is an extrapolation.
+Figure 9: (a) Modified Panda-gym benchmarks, (b) the training tasks, moderate test tasks, and extreme test tasks.
+The moderate test task setting involves combinatorial interpolation, while the extreme test task setting includes unseen ranges of environmental features and represents an extrapolation.
 D.1 Modified Panda-gym
-We modified the original Pick&Place task in Panda-gym [Gallouédec et al., 2021] by setting the $z$ dimension (i.e., the desired height) of the cube’s goal position equal to 00  [^3] If $z$ is not equal to 0, Pick&Place skill is always needed to solve tasks.
-and maintaining the freedom  [^4] In the original Push task, the grippers are blocked to ensure the agent can only push cubes.
-However, this hinders the agent from learning Pick&Place skills, leading to failure when it encounters an ”unpushable” scenario.
-of grippers, so the agent can explore whether it is supposed to push the cube or grasp it.
+We modified the original Pick&Place task in Panda-gym [Gallouédec et al., 2021] by setting the $z$ dimension (i.e., the desired height) of the cube’s goal position to 00  [^3] If $z$ is not equal to 0, the Pick&Place skill is always required to solve the tasks.
+and maintaining the freedom of the grippers  [^4] In the original Push task, the grippers are blocked to ensure the agent can only push cubes.
+However, this restriction prevents the agent from learning Pick&Place skills, leading to ”unpushable” failure in Figure 1 (b)., allowing the agent to explore whether it should push or grasp the cube.
 Skills in this benchmark are defined as:
-  • Pick&Place skill: This skill specifically refers to the agent trying to use the gripper to grasp the cube, pick it up off the table and place it in the goal position.
-We determine the Pick&Place skill by detecting no contact points between the table and the cube, two contact points between the robot’s end effector and the cube, and the cube’s height being greater than half its width, meaning it is on the table.
-  • Push skill: This skill refers to the agent giving a push force to the cube and then the cube slides to the goal position.
-We confirm the Push skill by detecting that the cube’s height is equal to half its width.
-  • Other skills: Besides the Pick&Place and Push skills, any other different modes of behaviour are classified as other skills.
-For example, with the Flip skill, the agent applies an initial force to the cube, causing it to roll on the surface or leave the table.
-In this case, we detect that the cube’s height is greater than half its width, and the number of contact points between the cube and the table is greater than zero.
-Some elements in the RL framework are defined as the following:
-State space: we use feature vectors which contain cube position (3 dimensions), cube rotation (3 dimensions), cube velocity (3 dimensions), cube angular velocity (3 dimensions), end-effector position (3 dimensions), end-effector velocity (3 dimensions), gripper width (1 dimension), desired goal (3 dimensions) and achieved goal (3 dimensions).
-Environmental features are not included in the state.
-Action space: the action space has 4 dimensions, the first three dimensions are the end-effector’s position changes and the last dimension is the gripper’s width change.
-During training, we randomly select a combination of environmental features from a training set sampling combinations from sets: $\text{mass}=1.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0\}$; $\text{mass}\in\{1.0,5.0,10.0\}$ and $\text{friction}=0.1$.
-At test time, we evaluate each algorithm in all tasks from the moderate test task setting, in which $\text{mass}\in\{5.0,10.0\}$ and $\text{friction}\in\{1.0,5.0,10.0\}$ (shown in Figure 8(b)), and all tasks from extreme test task setting: $\text{mass}=30.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0,30.0\}$; $\text{mass}\in\{1.0,5.0,10.0,30.0\}$ and $\text{friction}=30.0$ (shown in Figure 8(b)).
+  • Pick&Place skill: This skill specifically refers to the agent using the gripper to grasp the cube, lift it off the table, and place it in the goal position.
+We determine the Pick&Place skill by detecting no contact points between the table and the cube, two contact points between the robot’s end effector and the cube, and the cube’s height being greater than half its width.
+  • Push skill: This skill involves the agent moving the cube on the table to the goal position, either by dragging or sliding it.
+We confirm the Push skill by detecting that the cube’s height equals half its width.
+  • Other skills: Any behaviour modes other than Pick&Place and Push are classified as other skills.
+Some elements in the RL framework are defined as follows:
+State space: We use feature vectors that contain the cube’s position (3 dimensions), cube rotation (3 dimensions), cube velocity (3 dimensions), cube angular velocity (3 dimensions), end-effector position (3 dimensions), end-effector velocity (3 dimensions), gripper width (1 dimension), desired goal (3 dimensions), and achieved goal (3 dimensions).
+Environmental features are not included in states.
+Action space: The action space has 4 dimensions; the first three dimensions represent changes in the end-effector’s position, and the last dimension represents the change in the gripper’s width.
+During training, we randomly select a combination of environmental features from a training set by sampling combinations from the following sets: $\text{mass}=1.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0\}$; $\text{mass}\in\{1.0,5.0,10.0\}$ and $\text{friction}=0.1$.
+At test time, we evaluate each algorithm on all tasks from the moderate test setting, where $\text{mass}\in\{5.0,10.0\}$ and $\text{friction}\in\{1.0,5.0,10.0\}$ (shown in Figure 9(b)), and on all tasks from the extreme test setting: $\text{mass}=30.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0,30.0\}$; $\text{mass}\in\{1.0,5.0,10.0,30.0\}$ and $\text{friction}=30.0$ (shown in Figure 9(b)).
 D.2 Modified MuJoCo
-We follow exactly the same environmental settings as DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020].
-For Hopper, Half-cheetah, Ant, and SlimHumanoid, we use the environments from the MuJoCo physics engine and use implementation available from Clavera et al.
+Figure 10: Ten environments in modified MuJoCo benchmark.
+We extended the modified MuJoCo benchmark introduced in DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020].
+In our extension, four new environments were added (Walker, Crippled Hopper, Crippled Walker, HumanoidStandup), compared to the original benchmark.
+Additionally, in our experiments, we used a different task set design (Table 3) than those used in the DOMINO and CaDM papers.
+For Hopper, Walker, Half-Cheetah, Ant, HumanoidStandup, and SlimHumanoid, we used the MuJoCo physics engine environments and implemented settings from Clavera et al.
 [2019b] and Seo et al.
-[2020], and scale the mass of every rigid link by scale factor $m$, and scale damping of every joint by scale factor $d$.
-For Crippled Ant and Crippled Half-cheetah, we use implementation available from Seo et al.
-[2020] and scale the mass of every rigid link by scale factor $m$, scale damping of every joint by scale factor $d$, and randomly select one leg, and make it crippled to change the dynamic transitions.
-Figure 9: Six environments in modified MuJoCo benchmark.
-Generalisation performance is measured in two different regimes: moderate and extreme, where the moderate draws environment parameters from a closer range to the training range, compared to the extreme.
-We have provided our settings for training, extreme and moderate test tasks in Table 3.
-Table 3: Environment parameters used for MuJoCo benchmark.
+[2020], scaling the mass of every rigid link by a scale factor $m$ and the damping of every joint by a scale factor $d$.
+For Crippled Ant, Crippled Hopper, Crippled Walker, and Crippled Half-Cheetah, we used the implementation available from Seo et al.
+[2020], scaled the mass of each rigid link by a factor of $m$, scaled the damping of each joint by a factor of $d$, and randomly selected joints to be uncontrollable (i.e., masking the corresponding actions with 0).
+Generalisation performance is measured in two different regimes: moderate and extreme, where the moderate regime draws environmental features from a range closer to the training range compared to the extreme regime.
+Our settings for training, extreme, and moderate test tasks are provided in Table 3.
+Table 3: Environmental features used for MuJoCo benchmark.
 Training Test (Moderate) Test (Extrame) Episode Length
-Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 1000
-Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 1000
-Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 500
-SlimHumanoid $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80\}$ $d\in\{0.40,0.50,1.70,1.80\}$ 500
-Crippled Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Leg: $\{0,1,2\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Leg: $\{3\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Leg: $\{3\}$ 1000
-Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Leg: $\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Leg: $\{4,5\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Leg: $\{1\}$ 1000
+Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 1000
+Ant $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 1000
+Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 1000
+Crippled Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.0\}$ $d\in\{0.20,0.40,1.60,1.80,4.0\}$ Crippled Legs $R_{1}\in\{0,1,2\}$ 1000
+SlimHumanoid $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80\}$ $d\in\{0.40,0.50,1.70,1.80\}$ 1000
+HumanoidStandup $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80,4.00\}$ $d\in\{0.40,0.50,1.70,1.80,4.00\}$ 1000
+Walker $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 2000
+Crippled Walker $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ Crippled Joints (right leg) =$=\{0,1,2\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Joints (right leg) $\in$ $\{0,1,2\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ Crippled Joints (left leg) $\in$ $\{3,4,5\}$ 2000
+Crippled Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Legs $R_{1}=\varnothing$ or $R_{1}\in\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Legs $R_{1}\in\{0,1,2,3\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Legs $\{R_{1},R_{2}\}\in\{0,1,2,3,4,5\}$ $(R_{1}\neq R_{2})$ 2000
+Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Joints (front leg) $R_{1}\in$$\{3,4,5\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Joints (back leg) $R_{1}\in\{0,1,2\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Joints $\{R_{1},R_{2}\}\in\{0,1,2,3,4,5\}$ $(R_{1}\neq R_{2})$ 2000
 D.3 Implementation details
 In this section, we provide the implementation details for SaMI.
-Our codebase is built on top of the publicly released implementation Stable Baselines3 by Raffin et al.
-[2021] as well as the implementation of InfoNCE by Oord et al.
+We present the pseudo-code for using SaNCE during meta-training and meta-testing in Algorithms 1 and 2.
+Our codebase is built on top of the publicly released implementation of Stable Baselines3 by Raffin et al.
+[2021] and the implementation of InfoNCE by Oord et al.
 [2019].
-A public and open-source implementation of SaMI is available at https://github.com/uoe-agents/SaMI.
+A public, open-source implementation of SaMI is available at https://github.com/uoe-agents/SaMI.
 Base algorithm.
-We use SAC [Haarnoja et al., 2018] for downstream evaluation of the learned context embedding.
-SAC is an off-policy actor-critic method that uses the maximum entropy framework for soft policy iteration.
+We use SAC [Haarnoja et al., 2018] for the downstream evaluation of the learned context embedding.
+SAC is an off-policy actor-critic method that leverages the maximum entropy framework for soft policy iteration.
 At each iteration, SAC performs soft policy evaluation and improvement steps.
 We use the same SAC implementation across all baselines and other methods.
-We trained SAC agents for 1.6 million timesteps on Panda-gym and MuJoCo benchmarks (i.e.
-Half-cheetah, Ant, Crippled Half-cheetah, Crippled Ant, SlimHumanoid).
-We evaluated trained agents every $10,000$ timesteps on tasks with fixed random seeds.
-Table 4: Hyperparameters used in Panda-gym and MuJoCo benchmarks. Most hyperparameter values are unchanged across tasks except for contrastive batch size and SaNCE loss coefficient.
+During the meta-training phase, we trained agents for 1.6 million timesteps in each environment on the Panda-gym and MuJoCo benchmarks.
+For meta-testing, we evaluated 100 episodes in each environment, with tasks randomly sampled from the moderate and extreme task sets.
+Table 4: Hyperparameters used in the Panda-gym and MuJoCo benchmarks. Most hyperparameter values remain unchanged across tasks, except for the contrastive batch size and the SaNCE loss coefficient.
 Hyperparameter Value
@@
 Encoder architecture.
-For our method, the context encoder $\psi$ is modelled as a Long Short-Term Memory (LSTM) that produces a 128-dimensional hidden state vector, subsequently processed through a single-layer feed-forward network to generate a 6-dimensional context embedding.
-The actor and critic both use the same context encoder to embed trajectories.
+In our method, the context encoder $\psi$ is modelled as a Long Short-Term Memory (LSTM) network that produces a 128-dimensional hidden state vector, which is subsequently processed through a single-layer feed-forward network to generate a 6-dimensional context embedding.
+We aim for the agent to complete the three steps of "explore effectively, infer, adapt" within an episode.
+Therefore, we initialise the hidden state and cell state of the LSTM to zero at the start of each episode.
+Both the actor and critic use the same context encoder to embed trajectories.
 For contrastive learning, SaNCE utilises a momentum encoder $\psi^{*}$ to generate positive and negative context embeddings [Laskin et al., 2020, He et al., 2020].
-Formally, denoting the parameters of $\psi$ as $\theta_{\psi}$ and those of $\psi^{*}$ as $\theta_{\psi^{*}}$, we update $\theta_{\psi^{*}}$ by:
+Formally, denoting the parameters of $\psi$ as $\theta_{\psi}$ and those of $\psi^{*}$ as $\theta_{\psi^{*}}$, we update $\theta_{\psi^{*}}$ as follows:
 $\theta_{\psi^{*}}\leftarrow m\cdot\theta_{\psi}+(1-m)\cdot\theta_{\psi^{*}}.$ (13)
@@
 Only the parameters $\theta_{\psi}$ are updated by back-propagation.
-The momentum update in Equation 14 makes $\theta_{\psi^{*}}$ evolve more smoothly by having them slowly track the $\theta_{\psi}$ with $m\ll 1$ (e.g., $m=0.05$ in this research).
+The momentum update in Equation (13) makes $\theta_{\psi^{*}}$ evolve more smoothly by having them slowly track the $\theta_{\psi}$ with $m\ll 1$ (e.g., $m=0.05$ in this research).
 This means that the target values are constrained to change slowly, greatly improving the stability of learning.
@@
 Hardware. For each experiment run we use a single NVIDIA Volta V100 GPU with 32GB memory and a single CPU.
+Algorithm 1 SaNCE Meta-training
+0:  Batch of training tasks $\{e_{n}\}_{n=1,...,N}$ from $\xi_{train}(e)$, soft-update rate $m$;
+1:  Initialize RL replay buffer $\mathcal{B}_{RL}$, encoder replay buffer $\mathcal{B}_{enc}$;
+2:  Initialize parameters $\psi$ for context encoder, $\psi^{*}$ for momentum context encoder and $\phi$ for the off-policy SAC;
+3:  while not done do
+4:     for each task $e_{n}$ do
+5:        for Roll-out time steps do
+6:           for time step $t$ < maximum episode length $T$ do
+7:              Update context embedding $c_{n}\sim\psi(c_{n}|\tau_{c_{n},0:t})$
+8:              Roll-out policy $\pi_{c_{n}}(a_{t}|s_{t},c_{n})$ and accumulate transition $(s_{t},a_{t},r_{t},s_{t+1})$;
+9:           end for
+10:           Add trajectory $\tau_{c_{n}}=\{s_{0},a_{0},r_{0},s_{1},r_{1},...,s_{T},a_{T},r_{T}\}$ to replay buffer $\mathcal{B}_{RL}^{n}$ and $\mathcal{B}_{enc}^{n}$;
+11:        end for
+12:     end for
+13:     for each training step do
+14:        for each task $e_{n}$ do
+15:           Sample RL batch $\{\tau_{c_{n}}\}\sim\mathcal{B}_{RL}^{n}$;
+16:           Sample a positive sample $\tau_{c_{n}}^{+}$ for generating query with highest return, positive samples $\{\tau_{c_{n}}^{-}\}$ and negative samples $\{\tau_{c_{n}}^{-}\}$ for encoding positive and negative embeddings;
+17:           Update $\phi$ with RL loss $\mathcal{L}_{\text{RL}}$;
+18:           Update $\psi$ with SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and RL loss $\mathcal{L}_{\text{RL}}$;
+19:           $\theta_{\psi^{*}}\leftarrow m\cdot\theta_{\psi}+(1-m)\cdot\theta_{\psi^{*}}$;
+20:        end for
+21:     end for
+22:  end while
+Algorithm 2 SaNCE Meta-testing
+0:  Batch of training tasks $\{e_{n}\}_{n=1,...,N}$ from $\xi_{test}(e)$;
+1:  while not done do
+2:     for each task $e_{n}$ do
+3:        for each episode do
+4:           for time step $t$ < maximum episode length $T$ do
+5:              Update context embedding $c_{n}\sim\psi(c_{n}|\tau_{c_{n},0:t})$
+6:              Roll-out policy $\pi_{c_{n}}(a_{t}|s_{t},c_{n})$ and accumulate transition $(s_{t},a_{t},r_{t},s_{t+1})$;
+7:           end for
+8:        end for
+9:     end for
+10:  end while
 Appendix E Additional results
 E.1 Balance contrastive and RL updates: loss coefficient $\alpha$
-While past work has learned hyperparameters to balance the contrastive loss coefficient $\alpha$ relative to the RL objective [Jaderberg et al., 2016, Bachman et al., 2019], we use both the contrastive and RL objectives together with equal weight $\alpha=1.0$ to be optimal for the MuJoCo benchmark, and $\alpha=0.01$ for the Panda-gym benchmark, we also analyse the effect of loss coefficient $\alpha$ for CCM, SaTESAC and SaCCM in MuJoCo (Figure 11) and Panda-gym (Figure 10) benchmarks.
+While past work has optimised hyperparameters to balance the contrastive loss coefficient $\alpha$ relative to the RL objective [Jaderberg et al., 2016, Bachman et al., 2019], we use both the contrastive and RL objectives with equal weight, setting $\alpha=1.0$ for the MuJoCo benchmark and $\alpha=0.01$ for the Panda-gym benchmark.
+Additionally, we analyse the effect of the loss coefficient $\alpha$ for CCM, SaTESAC, and SaCCM in the MuJoCo (Figure 12) and Panda-gym (Figure 11) benchmarks.
 Figure 11: Loss coefficient $\alpha$ analysis of Panda-gym benchmark in training and test (moderate and extreme) tasks.
@@
 E.2.1 Buffer size
-Figure 12: Comparison of different buffer sizes in MuJoCo benchmark in training tasks (over 5 seeds). Buffer size = 400000, 100000, 10000, and 1000.
+Figure 13: Comparison of different buffer sizes in the MuJoCo benchmark on training tasks (averaged over 5 seeds). Buffer sizes are 400,000, 100,000, 10,000, and 1,000.
 E.2.2 Contrastive batch size
-Figure 13: Comparison of different contrastive batch sizes in MuJoCo benchmark in training tasks (over 5 seeds). contrastive batch size = 512, 128, 16, and 8.
+Figure 14: Comparison of different contrastive batch sizes in the MuJoCo benchmark on training tasks (averaged over 5 seeds). Contrastive batch sizes are 512, 128, 16, and 8.
 Appendix F Further skill analysis
@@
 F.1.1 Visualisation of context embedding
-We visualise the context embedding via t-SNE [Van der Maaten and Hinton, 2008] (Figure 15) and PCA [Jolliffe, 2002] (Figure 14).
-We found that when the mass of the cube is high (30 Kg and 10 Kg), the agent learned the Push skill (the yellow bounding box in Figure 5.1(a)), whereas with lower masses, the agent learned the Pick&Place skill.
+We visualise the context embedding using UMAP [McInnes et al., 2020] (Figure LABEL:fig:vis_umap_panda) and t-SNE [Van der Maaten and Hinton, 2008] (Figure 16).
+When the mass of the cube is high (30 kg and 10 kg), the agent learned the Push skill (indicated by the yellow bounding box in Figure 5.1(a)), whereas with lower masses, the agent learned the Pick&Place skill.
 However, as shown in Figure 16(b), CCM did not display clear skill grouping.
-This indicates that SaMI extracts high-quality skill-related information from the trajectories and helps with embodying diverse skills.
-Figure 14: PCA visualization of context embedding extracted from trajectories collected in Panda-gym environments.
-Figure 15: t-SNE visualization of context embedding extracted from trajectories collected in Panda-gym environments.
+This indicates that SaMI extracts high-quality skill-related information from the trajectories and enables agents to autonomously discover a diverse range of skills for handling multiple tasks.
+Figure 15: UMAP visualisation of context embeddings extracted from trajectories collected in the Panda-gym environments.
+Figure 16: t-SNE visualisation of context embeddings extracted from trajectories collected in the Panda-gym environments.
 F.1.2 Heatmap of Panda-gym benchmark
-Figure 16: Heatmap of success rate and learned skills of SaCCM.
-For each grid, the $(i,j)^{th}$ location shows the probability of the skills executed in 100 evaluations with $(\text{mass}=i,\text{friction}=j)$.
-We determine the skills by detecting contact points between the end effector and the cube and between the cube and the table.
-This section provides the heatmap results and further analysis of TESAC, CCM, and SaTESAC on the Panda-gym benchmark.
-Initially, from the heatmap results of SaTESAC and SaCCM (Figure 6), we observed that with higher cube masses (30 and 10 kg), the agent adopted the Push skill (as indicated by the clustered points in the yellow bounding box in Figure 5.1(a)).
-At lower masses, the agent adopted the Pick&Place skill.
-In contrast to CCM, as seen in Figures 16(e-f), CCM predominantly learned the Pick&Place skill, resulting in a decline in success rates for tasks at $mass=30$, as the agent could not lift the cube off the table using the Pick&Place skill, as depicted in Figure 16(d).
-The visualisation of the context embedding (Figure 15) did not show clear grouping across different tasks.
-Finally, TESAC only mastered the Push skill.
-The Push skill is relatively simpler to learn than the Pick&Place skill, as it does not require the agent to master manipulating fingers to pick up cubes.
+Figure 17: Heatmap of success rates and learned skills of SaCCM.
+For each grid, the $(i,j)^{\text{th}}$ location shows the probability of the skills executed over 100 evaluations with $(\text{mass}=i,\text{friction}=j)$.
+This section presents the heatmap results and further analysis of TESAC, CCM, and SaTESAC on the Panda-gym benchmark.
+From the heatmap results of SaTESAC and SaCCM (Figure 6), we observe that, with higher cube masses (30 and 10 kg), the agent executed the Push skill (indicated by the clustered points within the yellow bounding box in Figure 5.1(a)).
+At lower masses, the agent executed the Pick&Place skill.
+In contrast, as shown in Figures 17(e-f), CCM predominantly learned the Pick&Place skill, resulting in a drop in success rates for tasks with $mass=30$, as the agent could not lift the cube off the table using the Pick&Place skill, as depicted in Figure 17(d).
+The visualisation of the context embedding (Figure 16) did not reveal clear skill grouping across different tasks.
+Finally, TESAC primarily mastered the Push skill.
+The Push skill is relatively simpler to learn than the Pick&Place skill, as it does not require the agent to manipulate its fingers to pick up cubes.
 Consequently, TESAC’s success rate notably decreased in environments with higher friction.
 F.2 MuJoCo
-We find that SaMI help the RL agents be versatile and embody multiple skills.
-For example, Figure 17 displays the t-SNE visualization results of four methods in the Crippled Half-cheetah environment for both test (moderate and extreme) and training tasks.
-Tasks corresponding to different crippled joints on the same leg are clustered together.
-Further, we have displayed rendered videos (https://github.com/uoe-agents/SaMI) to better demonstrate the different skills learned.
-Figure 17: t-SNE and PCA visualisation of context embedding extracted from trajectories collected in Crippled Half-cheetah environment.
-Figure 18: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Crippled Ant environment.
-Figure 19: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Half-cheetah environment.
-Figure 20: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Ant environment.
-Figure 21: t-SNE and PCA visualization of context embedding extracted from trajectories collected in SlimHumanoid environment.
-Figure 22: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Hopper environment.
-Table 5: Sub-standard comparison of average return with baselines in MuJoCo benchmark (over 5 seeds).
-The blue background indicates that our method surpasses the baseline, and bold text signifies the highest average return.
+SaMI enables RL agents to be versatile and embody multiple skills. Additionally, video demos (available at https://github.com/uoe-agents/SaMI) provide a better demonstration of the different skills.
+Figure 18: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in the Crippled Half-Cheetah environment.
+"cripple_set" refers to the index of the crippled joint.
+The figure shows the context embeddings of tasks in three moderate test settings.
+Combined with the video demos2 for skill analysis, the Crippled Half-Cheetah robot executed three distinct forward running skills.
+Figure 19: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in the Crippled Ant environment.
+"cripple_set" refers to the index of the crippled joint.
+When 3 or 4 legs are available, the Ant robot (trained with SaCCM) rolls to adapt to varying mass and damping.
+However, with only 2 adjacent legs during zero-shot generalisation, it switches to walking.
+If 2 opposite legs are available, the Ant can still roll but eventually tips over.
+Please refer to the video demos2.
+Figure 20: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in the Half-Cheetah environment.
+During zero-shot generalisation, SaCCM demonstrates different skills for running forwards at various speeds, as well as skills for doing flips and faceplanting.
+Figure 21: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Ant environment.
+The Ant robot learned a single skill, rolling, to adapt to different mass and damping values.
+Figure 22: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in SlimHumanoid environment.
+The Humanoid Robot crawls on the ground using one elbow.
+When the damping is relatively high (damping=1.8), the Humanoid Robot can crawl forward stably, but when the damping is low (damping=0.4), it tends to roll.
+Figure 23: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in HumanoidStandup environment.
+SaCCM and SaTESAC learned a sitting posture that makes it easier to stand up, allowing it to generalise well when mass and damping change.
+Figure 24: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Hopper environment.
+Combined with the video demos 2 for skill analysis, the plots for SaCCM and SaTESAC show two skills: 1) when the mass is low, the Hopper hops in an upright posture; 2) when the mass is higher, the Hopper hops forward on the floor.
+Figure 25: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Crippled Hopper environment.
+"cripple_set" refers to the index of the crippled joint.
+SaCCM and SaTESAC learned to take a big hop forward at the beginning (i.e., effective exploration) and then switch to different skills based on environmental feedback.
+Figure 26: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Walker environment.
+The Walker learned a single skill, hopping forward on the floor through both right and left legs.
+Figure 27: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Crippled Walker environment.
+"cripple_set" refers to the index of the crippled joint.
+The Crippled Walker (trained with SaTESAC and SaCCM) learned to hop forward using the right leg (cripple_set:0 and cripple_set:2) in the training and moderate test tasks and switched to hopping forward using the left leg (cripple_set:4) in the extreme test tasks.
+Table 5: Environmental features used for MuJoCo benchmark from DOMINO and CaDM.
+Training Test (Moderate) Test (Extrame) Episode Length
+Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 1000
+Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 1000
+SlimHumanoid $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80\}$ $d\in\{0.40,0.50,1.70,1.80\}$ 500
+Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Joints: $\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Joints: $\{4,5\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Joints: $\{4,5\}$ 1000
+Table 6: Comparison of average return $\pm$ standard deviation with baselines in MuJoCo benchmark (over 5 seeds).
+The bold text signifies the highest average return.
 The numerical results for PPO+DOMINO are copied from Mu et al.
@@
 PE-TS+CaDM 2848.4$\pm$61.9 2121.0$\pm$60.4 1200.7$\pm$21.8 8264.0$\pm$1374.0 7087.2$\pm$1495.6 4661.8$\pm$783.9
-SaTESAC 911$\pm$54 590$\pm$111 552$\pm$119 7303$\pm$1353 4077$\pm$424 1723$\pm$88
-SaCCM 1015$\pm$60 629$\pm$126 588$\pm$92 7662$\pm$316 4384$\pm$223 1828$\pm$52
-SlimHumanoid Handcapped Half-Cheetah
+SaTESAC 908$\pm$65 640$\pm$117 532$\pm$88 7430$\pm$1026 4058$\pm$890 1780$\pm$102
+SaCCM 928$\pm$141 635$\pm$94 555$\pm$88 7154$\pm$965 3849$\pm$689 1926$\pm$218
+SlimHumanoid Crippled Half-Cheetah
 Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
@@
 PE-TS+CaDM 1371.9$\pm$400.0 903.7$\pm$343.9 814.5$\pm$274.8 3294.9$\pm$733.9 2618.7$\pm$647.1 1294.2$\pm$214.9
-SaTESAC 10501$\pm$2102 8483$\pm$1502 6669$\pm$883 5380$\pm$917 2334$\pm$169 1848$\pm$341
-SaCCM 9738$\pm$92 8389$\pm$231 5606$\pm$2027 6210$\pm$402 2656$\pm$231 2070$\pm$651
+SaTESAC 10216$\pm$1620 7886$\pm$2203 6123$\pm$1403 5169$\pm$730 2184$\pm$592 1628$\pm$281
+SaCCM 9312$\pm$705 7430$\pm$1587 6473$\pm$2001 5709$\pm$744 2795$\pm$446 2115$\pm$466
 Appendix G A comparison with DOMINO and CaDM
-In this section, we give a brief comparison between our methods and methods from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] in the MuJoCo benchmark because we are using the exact same environmental setting.
+In this section, we give a brief comparison between our methods and methods from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] in the MuJoCo benchmark because we are using the exact same environmental setting (shown in Table 5).
 DOMINO [Mu et al., 2022] is based on the InfoNCE $K$-sample estimator.
@@
 In contrast, SaTESAC and SaCCM, trained for 1.6 million timesteps without pre-trained encoders, achieve considerably higher average returns across four environments (Table 6).
-Therefore, it is crucial to focus on extracting mutual information in contrastive learning that directly optimises downstream tasks, integrating rather than segregating representation learning from task performance.
+Therefore, it is crucial to focus on extracting MI in contrastive learning that directly optimises downstream tasks, integrating rather than segregating representation learning from task performance.
 CaDM [Lee et al., 2020] proposes a context-aware dynamics model adaptable to changes in dynamics.
@@
 An insight here is that our method outperforms the model-free CaCM approach, but not the model-based one.
+This is consistent with what is empirically found in CaDM [Lee et al., 2020]: prediction models are more effective when the transition function changes across tasks.
 Therefore, we consider that a model-based approach to SaMI could be an interesting extension for future work.
+Appendix H Statistical hypothesis tests (paried t-tests)
+We used a t-test [Rice and Rice, 2007] to conduct a statistical hypothesis test to determine whether SaMI brought a statistically significant improvement.
+we reported the p-value of the t-test in MuJoCo (Table 7) and Panda-gym (Table 8) benchmarks.
+* next to the number is used to indicate that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI at a significance level of 0.05.
+From Table 7 and 8, SaMI brings significant improvement on the extreme test set in which the RL agent needs to execute diverse skills.
+The statistically significant test aligns with our results in the skill analysis (i.e., video demos, t-SNE and UMAP visualisation).
+In complex environments that require high skill diversity from the RL agent, the statistically significant improvement and higher returns/success rates brought by SaMI are evident.
+Table 7: The p-value of the statistical hypothesis tests (paried t-tests) for comparing the effectiveness of SaMI in MuJoCo benchmark (over 5 seeds).
+* next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI at a significance level of 0.05.
+The “SaTESAC-TESAC” row indicates the p-value for the return improvement brought by SaMI to TESAC; the “SaCCM-CCM” row indicates the p-value for the return improvement brought by SaMI to CCM.
+Crippled Ant Crippled Half-cheetah
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+SaTESAC-TESAC 0.121 0.109 9.54E-07 0.154 0.0024 0.0889
+SaCCM-CCM 0.913 0.108 0.008 0.04 0.307 0.106
+Ant Half-cheetah
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+SaTESAC-TESAC 0.136 0.034 0.021 0.346 0.224 0.004
+SaCCM-CCM 0.138 0.275 0.163 0.73 0.791 0.005
+SlimHumanoid HumanoidStandup
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+SaTESAC-TESAC 0.139 0.456 0.006 0.037 0.129 0.003
+SaCCM-CCM 0.113 0.059 0.008 0.048 0.027 0.01
+Hopper Crippled Hopper
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+SaTESAC-TESAC 0.747 0.089 0.707 0.459 0.69 0.088
+SaCCM-CCM 0.52 0.599 0.969 0.967 0.897 0.0002
+Walker Crippled Walker
+Training Test (moderate) Test (extreme) Training Test (moderate) Test (extreme)
+SaTESAC-TESAC 0.312 0.082 0.003 0.223 0.048 0.028
+SaCCM-CCM 0.55 0.541 0.022 0.794 0.079 0.011
+Table 8: The p-value of the statistical hypothesis tests (paried t-tests) for comparing the effectiveness of SaMI in Panda-gym benchmark (over 5 seeds).
+* next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI at a significance level of 0.05.
+The “SaTESAC-TESAC” row indicates the p-value for the return improvement brought by SaMI to TESAC; the “SaCCM-CCM” row indicates the p-value for the return improvement brought by SaMI to CCM.
+Training Test (moderate) Test (extreme)
+SaTESAC-TESAC 0.000260 0.000160 0.001310
+SaCCM-CCM 0.000230 0.002190 0.001390